### v1

# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

Abstract
We propose a new asymptotic equipartition property for the perplexity of a large piece of text generated by a language model and present theoretical arguments for this property. Perplexity, defined as a inverse likelihood function, is widely used as a performance metric for training language models. Our main result states that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions. This means that language models are constrained to only produce outputs from a “typical set”, which we show, is a vanishingly small subset of all possible grammatically correct outputs. We present preliminary experimental results from an open-source language model to support our theoretical claims. This work has possible practical applications for understanding and improving “AI detection” tools and theoretical implications for the uniqueness, predictability and creative potential of generative models.
1 Introduction and Motivation.
Consider a generative model, defined as an algorithm that takes a user input ${\mathbf{X}}$ and produces an output ${\mathbf{Y}}$ that statistically resembles data from a natural source. A specific type of generative model is a large language model (LLM) whose output ${\mathbf{Y}}$ is a body of text and the input is a text user prompt ${\mathbf{X}}$. State-of-the-art LLMs (Anthropic, 2024; OpenAI, 2024) are now able to produce detailed and information-rich text outputs such as entire screenplays and book-length manuscripts from a short and simple user prompt. Furthermore, LLMs produce text that can imitate human language well-enough to pass the Turing test (Pinar Saygin et al., 2000) i.e. resembles text created by humans well enough to be convincing to human observers. In this work, we argue that an LLM is strongly constrained by statistical laws. Specifically, we show that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions. This means that any language model is constrained to only output text strings from a typical set, which is an exponentially vanishing subset of all possible grammatically correct strings.
1.1 Contribution: Equipartition Property for Perplexity
Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister and Cotterell, 2021). It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover and Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg and Keller, 2008; Cohen and Pakhomov, 2020) phenomena at least partially. Many “AI detection” tools for identifying synthetic text from language models are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory. The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn.
The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara and Morita, 2000; Timo et al., 2010). The outputs of non-trivial language models, however, cannot be reasonably modeled as a sequence of nearly-independent symbols (Brown et al., 2020).
2 Background and Definitions
Let M be a generative model described by ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$, where the output ${\mathbf{Y}}$ consisting of a (potentially infinite) string of tokens ${\mathbf{Y}}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}~{}\dots]$ is a deterministic function of user prompt ${\mathbf{X}}$ and a pseudo-random sequence ${\mathbf{W}}$. Each token is chosen from a finite set of tokens $Y_{n}\in\mathcal{T}$. The model M can produce a number of different output strings ${\mathbf{Y}}$ for the same user prompt ${\mathbf{X}}$ corresponding to different values of the pseudo-random sequence ${\mathbf{W}}$. This defines a probability distribution $P({\mathbf{Y}}|{\mathbf{X}})$, or more formally, a sequence of probability distributions $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ where ${\mathbf{Y}}_{N}\doteq[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ is the substring of ${\mathbf{Y}}$ consisting of the first $N$ tokens.
2.1 Sequential LLM
Practical implementations of LLMs specify the probability distribution iteratively (Radford et al., 2018):
$P(Y_{1},Y_{2}|{\mathbf{X}})$ $=P(Y_{1}|{\mathbf{X}})P(Y_{2}|Y_{1},{\mathbf{X}})$ (1)
and so on. Thus, the model M first draws a random value for the first token say $Y_{1}=y_{1}$ by sampling from the distribution $P(Y_{1}|{\mathbf{X}})$. Then the model determines a distribution for the second token as a function of the initial prompt ${\mathbf{X}}$ and the randomly chosen first token $y_{1}$. Thus, the second token is randomly sampled from a distribution $P(Y_{2}|{\mathbf{X}},Y_{1}=y_{1})$ and so on. We can write
$P({\mathbf{Y}}_{N}|{\mathbf{X}})$ $=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
Given a string ${\mathbf{Y}}$, open-source LLMs can be programmed to print out the distributions $p_{n}(Y_{n})$ from which its tokens were selected. Specifically, given a user prompt ${\mathbf{X}}$ and a string of tokens ${\mathbf{Y}}_{N}\equiv[Y_{1},~{}Y_{2},~{}\dots,~{}Y_{N}]$, it is possible to get a complete listing of the distributions $p_{n}(Y_{n}),~{}n=1\dots N$. Note that complete knowledge of the $p_{n}(y)$ is not the same as complete knowledge of $P({\mathbf{Y}}|{\mathbf{X}})$ even for a fixed user prompt ${\mathbf{X}}$. As an example, the former requires knowledge of $p_{3}(y)\equiv P(y|Y_{1},Y_{2},{\mathbf{X}})$ for different values of $y\in\mathcal{T}$, but only conditioned on the specific values of $Y_{1},~{}Y_{2}$ contained in one specific string ${\mathbf{Y}}$.
Remark. Equation (2) is simply an application of the Bayes rule of probability theory and it always holds for any generative model regardless of whether the tokens $Y_{n}$ are sequentially generated. However, the conditional distributions $p_{n}(y)$ are not in general easily accessible, so while (2) is true for all generative models, it may only be useful for sequential models.
The perplexity $p_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$. It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(p_{M}({\mathbf{Y}}_{N})\right)\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
2.2 A Toy Problem
Let $\alpha(y),~{}\beta(y)$ be two fixed probability distributions over the (discrete) set $\mathcal{T}$ of tokens. Consider a toy problem involving two language models A and B, that each generate a string of tokens ${\mathbf{Y}}=[Y_{1},~{}Y_{2},~{}\dots]$ where each token is generated iid from the distribution $\alpha(y)$ and $\beta(y)$ respectively. The iid assumption implies that the tokens $Y_{n}$ can be thought of as being generated by a stationary and ergodic random process [^1] Strictly speaking, a stationary random process “starts” at $n=-\infty$ rather than $n=1$; here we ignore the tokens produced by the assumed random process before $n=1$..
Consider a long string ${\mathbf{A}}_{N}\doteq[A_{1},~{}A_{2},~{}\dots,~{}A_{N}],~{}N\gg 1$ randomly generated from model A. Let $p_{\mathbf{A}}(y)$ denote the empirical distribution of $y\in\mathcal{T}$ in the string ${\mathbf{A}}_{N}$:
$p_{\mathbf{A}}(y)$ $\doteq\frac{n_{\mathbf{A}}(y)}{N}\equiv\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}(A_{n}\equiv y),~{}\forall y\in\mathcal{T}$ (4)
where $\mathbf{1}(.)$ denotes the indicator function and $n_{\mathbf{A}}(y)$ is the relative frequency of token $y$ in the (long) string ${\mathbf{A}}_{N}$. The log-perplexity of string ${\mathbf{A}}_{N}$ for model A is:
$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
where $H(\gamma_{1},\gamma_{2})\doteq-\sum_{y\in\mathcal{T}}\gamma_{1}(y)\log_{2}(\gamma_{2}(y))$ is the cross-entropy between two distributions $\gamma_{1}(y),~{}\gamma_{2}(y)$ over $y\in\mathcal{T}$.
It is well-known $H(\gamma_{1},\gamma_{2})\geq H(\gamma_{1})$ with equality when $\gamma_{1}\equiv\gamma_{2}$, where $H(\gamma)\doteq-\sum_{y\in\mathcal{T}}\gamma(y)\log_{2}(\gamma(y))$ is the entropy of distribution $\gamma$. Thus we have from (5):
$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(\alpha)~{}{\mathrm{with~{}equality~{}iff~{}}}p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
The simplest version of the classical Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory states that the log-perplexity of a long string ${\mathbf{A}}_{N}$ of iid symbols is almost always very close to the entropy $H(\alpha)$ of the distribution $\alpha(y)$ the symbols are drawn from.
Proposition 1.
Simple AEP. For a long text string ${\mathbf{A}}_{N}=[A_{1},~{}A_{2},~{}\dots,~{}A_{N}]$ of iid tokens $A_{n}$ drawn from a distribution $\alpha(y)$, the log perplexity $l_{A}({\mathbf{A}}_{N})$ as defined in (3) is close to the entropy of the distribution $H(\alpha)$ of the distribution $\alpha(y)$ with high probability:
$\lim_{N\rightarrow\infty}\Pr[|l_{A}({\mathbf{A}}_{N})-H(\alpha)|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (7)
Proposition 1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e. convergence in probability. applied to the random variable $l_{M}({\mathbf{A}}_{N})$:
$l_{M}({\mathbf{A}}_{N})$ $\rightarrow E\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
From (5) and (6), we see that (8) itself is equivalent to $p_{\mathbf{A}}(y)\equiv\frac{n_{\mathbf{A}}(y)}{N}\rightarrow\alpha(y),~{}\forall y\in\mathcal{T}$ i.e. the empirical distribution $p_{\mathbf{A}}(y)$ for long strings ${\mathbf{A}}$ is very close to $\alpha(y)$ with high probability. Putting these observations together we have:
$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(\alpha)\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
Intuitively, (9) states that a long string ${\mathbf{A}}$ generated from model A is likely to have a lower perplexity for model A than for any other model B. This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts. This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided that we accept the Shannon model of language i.e. the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961).
2.3 A Naive Method for Detecting Text Generated by a Language Model
Given a text string ${\mathbf{Y}}$, we seek to determine whether or not it was generated by a given model M. This is the problem of “white-box” detection of AI generated text where we assume knowledge of the specific language model that potentially generated the text. A naive method for doing this is as follows: compare the log-perplexity $l_{M}({\mathbf{Y}})$ of the text ${\mathbf{Y}}$ for model M with a threshold: if $l_{M}({\mathbf{Y}})<l_{th}$, we determine that ${\mathbf{Y}}$ is likely to have been generated by model M.
Consider a long string ${\mathbf{B}}\doteq[B_{1},~{}B_{2},~{}\dots,~{}B_{N}]$ randomly generated from model B. Using the same argument as in Section 2.2, we have:
$l_{B}({\mathbf{B}})$ $\equiv H(p_{\mathbf{B}},\beta)\rightarrow H(\beta)\leq H(p_{\mathbf{B}},\alpha)\equiv l_{A}({\mathbf{B}})$ (10)
Note, however, that we have no basis to assert any relationship between $l_{A}({\mathbf{A}})$ and $l_{A}({\mathbf{B}})$. Indeed if $H(\beta)\ll H(\alpha)$ i.e. the distribution $\beta$ has significantly lower entropy than $\alpha$, it cannot be ruled out that $l_{A}({\mathbf{B}})<l_{A}({\mathbf{A}})$ i.e. a text-string generated by a different model has lower perplexity for a model than a text-string generated by itself.
The mathematical properties of cross-entropy provide a rigorous basis for using perplexity to compare the compatibility of different models with a given text string. However, this theory does not support using perplexity to compare different text strings for compatibility with a given model.
2.4 A Modest Generalization
This theoretical analysis relies on Proposition 1 which only applies to toy models that generate strings of iid tokens. Of course, the simple AEP in Proposition 1 has been extended in various ways in the literature such as the following version (see Tal et al. (2017), Appendix B.4), which we will make use of in the sequel.
Proposition 2.
Generalized AEP. Consider a language model A’ that generates text string ${\mathbf{A}}^{\prime}_{N}=[A^{\prime}_{1},~{}A^{\prime}_{2},~{}\dots,~{}A^{\prime}_{N}]$ where the tokens $A^{\prime}_{n}$ are drawn independently from a sequence of distributions $\alpha_{n}(y),~{}n=1,2,\dots$. Assuming the entropies of distributions $\alpha_{n}(y)$ are uniformly bounded i.e. $H(\alpha_{n}(y))<H_{max}<\infty,~{}\forall n$, the log perplexity $l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})$ of the string ${\mathbf{A}}^{\prime}_{N}$ as defined in (3) is close to the average entropy of the distributions $\alpha_{n}(y),~{}n=1,2,\dots,N$ with high probability:
$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,~{}\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (11)
The regularity condition that the entropies of $\alpha_{n}(y)$ are uniformly bounded is trivially true if the token dictionary $\mathcal{T}$ is a finite set: $H(p(y))\leq\log_{2}|\mathcal{T}|$ for any distribution $p(y)$ over $y\in\mathcal{T}$.
While Proposition 2 does not lend itself to an intuitive interpretation in terms of the empirical distribution of the tokens $A^{\prime}_{n}$, it too is a direct consequence of the Law of Large Numbers applied to the log-perplexity random variable. Much of the analysis in Section 2.2 can be extended to generalized models like A’ in Proposition 2 that allows each token $A^{\prime}_{n}$ to be drawn from different distributions $\alpha_{n}(y)$. However, the tokens in these models must be drawn from fixed distributions independent of past tokens. This is still quite trivial compared to modern language models where the output tokens depend in highly complex and sophisticated ways on past tokens.
3 An Equipartition Property for Perplexity
We now propose a generalization of the theory described in Section 2.2 to a more interesting class of models. Consider a language model M and a random infinitely long text string ${\mathbf{Y}}$ generated by M, whose probabilities for a given prompt ${\mathbf{X}}$ is described by (2). Specifically, given a model M and a string ${\mathbf{Y}}$, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{T}$. We will assume that the prompt ${\mathbf{X}}$ is fixed.
Definition 1.
The empirical entropy $h_{M}({\mathbf{Y}}_{N})$ of model M for string ${\mathbf{Y}}_{N}$ is defined as:
$h_{M}({\mathbf{Y}}_{N})\doteq\frac{1}{N}\sum_{n=1}^{M}H(p_{n})$ (12)
where $H(p)\equiv-\sum_{y\in\mathcal{T}}p(y)\log_{2}(p(y))$ is the entropy of the distribution $p(y)$ over $y\in\mathcal{T}$.
Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number. Indeed, we can show that
$E\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv E\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$ (13)
where the expectation is over random text strings ${\mathbf{Y}}_{N}$.
Definition 2.
Let the log-deviation $\lambda(p)$ of a probability distribution $p(y)$ over $y\in\mathcal{T}$ be defined as:
$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}}~{}\mathrm{where}~{}S(p)\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (14)
Note that the entropy $H(p)$ and log-deviation $\lambda(p)$ are the mean and standard deviation of the log-likelihood random variable $l_{p}(y)\equiv-\log_{2}p(y)$ under distribution $p(y)$.
Definition 3.
The log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of a string ${\mathbf{Y}}_{N}$ for model $M$ is defined as:
$\lambda_{M}({\mathbf{Y}}_{N})$ $\doteq\frac{1}{N}\sqrt{\sum_{n=1}^{N}\lambda^{2}(p_{n})}$ (15)
We can again interpret $H(p_{n}),~{}\lambda(p_{n})$ as the conditional mean and standard deviation of $-\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$. Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string ${\mathbf{Y}}$ are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes.
Preview of our main result. We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$. But this is very wrong - while the log-deviation $\lambda(p)$ of a fixed probability distribution $p(y)$ is a deterministic number, $\lambda_{M}({\mathbf{Y}}_{N})$ of a string ${\mathbf{Y}}_{N}$ is a random variable that takes different numerical values for different random text strings ${\mathbf{Y}}$. This considerably complicates our analysis. Fundamentally, their randomness arises from the fact that the distributions $p_{n}(y)$ are different for different strings ${\mathbf{Y}}$. This is because the tokens $Y_{n}$ in the string ${\mathbf{Y}}$ are not generated independently; instead future tokens strongly depend on earlier tokens. This is an essential property of modern language models that allows them to generate sophisticated text outputs. We now introduce the conceptual device of an auxiliary language model to get around this difficulty.
3.1 Concept of an Auxiliary Language Model
Let us consider a fixed string ${\mathbf{Y}}$. For this particular string ${\mathbf{Y}}$ and model M, we have a fixed set of distributions $p_{n}(y)$. Now define an auxiliary generative model $M^{\prime}_{\mathbf{Y}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~{}Y_{2}^{\prime},~{}\dots,~{}Y_{N}^{\prime},~{}\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $p_{n}(y)$. This construct of the auxiliary model $M^{\prime}_{\mathbf{Y}}$ for a fixed string ${\mathbf{Y}}$ allows us to work with fixed distributions $p_{n}(y)$ which are not themselves functions of a random string.
In particular, note that the empirical entropy $h_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ and log-deviation $\lambda_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ for model $M^{\prime}_{\mathbf{Y}}$ are both deterministic functions of $N$ independent of the string ${\mathbf{Y}}^{\prime}_{N}$. Specifically, we have $h_{M^{\prime}_{\mathbf{Y}}}(N)\equiv h_{M}({\mathbf{Y}}_{N})$. Similarly, we have $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)\equiv\lambda_{M}({\mathbf{Y}}_{N})$ which, since we have fixed our special string ${\mathbf{Y}}$, is now also fixed. Indeed, it is easy to show that $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)$ is the standard-deviation of the random variable $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ the log-likelihood of the random string ${\mathbf{Y}}^{\prime}_{N}$. Then, by the Chebyshev Inequality (Cohen, 2015), we have:
$\Pr\left(\left|l_{M}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{\mathbf{Y}}}(N)\right|>\epsilon\right)$ $\leq\frac{\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)}{\epsilon^{2}}$ (16)
Clearly, for substrings ${\mathbf{Y}}_{N}$ of our special string ${\mathbf{Y}}$ that defines the auxiliary model $M^{\prime}_{\mathbf{Y}}$, we have by definition $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N})$. In addition, the entropy $H(P({\mathbf{Y}}^{\prime}_{N}))$ of the random strings ${\mathbf{Y}}^{\prime}_{N}$ output by auxiliary model $M^{\prime}_{\mathbf{Y}}$ is by definition $H(P({\mathbf{Y}}^{\prime}_{N}))\equiv h_{M}({\mathbf{Y}}_{N})$.
The auxiliary model $M^{\prime}_{\mathbf{Y}}$ is much simpler than the general language model M. In particular, because the tokens $Y^{\prime}_{n}$ are generated independently of other tokens, the AEP in Proposition 2 applies to strings ${\mathbf{Y}}^{\prime}$ generated by the auxiliary model $M^{\prime}_{\mathbf{Y}}$. Indeed we note from (15) that the log-deviation $\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)$ vanishes for large $N$ under some regularity conditions on the distributions $p_{n}(y)$, in which case, the AEP for model $M^{\prime}_{\mathbf{Y}}$ follows from (16).
3.2 Ensemble of Auxiliary Models
We can think of a generative model M as a statistical ensemble of a large number of (simpler) auxiliary models. Each auxiliary model is capable of generating a random text string consisting of a sequence of independent tokens. We can then imagine choosing randomly from one of the auxiliary models. We now make this idea more precise.
Consider an enumeration ${\mathbf{Y}}^{(1)},~{}{\mathbf{Y}}^{(2)},~{}\dots,~{}\dots,~{}{\mathbf{Y}}^{(N|\mathcal{T}|)}$ of the possible text strings ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$. Let $M_{{\mathbf{Y}}^{(n)}}$ denote the auxiliary model associated with model M and string ${\mathbf{Y}}^{(n)}$. A sequential generative model can be represented by a probability tree where each branch in the tree represents a new token being generated. The model M as well as each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can each be represented by such a probability tree. An example of a probability tree representing a generative model and a sample random string ${\mathbf{Y}}$ represented by a sequence of branches through the tree are shown in Fig. 1 which also shows the probability tree for one possible auxiliary model $M^{\prime}_{\mathbf{Y}}$.
Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens. Also shown in red is the probability tree of an auxiliary model for a string ${\mathbf{Y}}$ with $Y_{1}=3,~{}Y_{2}=1$ and so on.
A generative model uses a pseudo-random number generator (PRNG) that produces a sequence of random symbols ${\mathbf{W}}$ to randomly choose a sequence of branches through its probability tree. As an example, suppose ${\mathbf{W}}=[W_{1},~{}W_{2},~{}\dots W_{N}~{}\dots]$ where the $W_{n}$ are independent and uniformly distributed in $[0,1]$ and a generative model uses $W_{n}$ to generate the n’th token $Y_{n}$ according to the distribution $p_{n}(y)$ using a sampling algorithm such as the classical inversion method Martino et al. (2018). Once we fix the PRNG and the sampling method, we can in principle write down a deterministic function to describe the generative model M as ${\mathbf{Y}}_{N}=m({\mathbf{X}},{\mathbf{W}}_{N})$ where recall that ${\mathbf{X}}$ is the fixed initial prompt. Each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can also similarly be described by deterministic functions ${\mathbf{Y}}_{N}=m_{{\mathbf{Y}}^{(n)}}({\mathbf{X}},{\mathbf{W}}_{N})$.
We can then define a new generative model that we call the ensemble model $M_{ensemble}$ as follows. Pick a pseudo-random sequence ${\mathbf{W}}_{1}$ and generate a string ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$ from model M. Now consider the auxiliary model $M_{{\mathbf{Y}}_{1}}$ corresponding to string ${\mathbf{Y}}_{1}$. Generate a pseudo-random string ${\mathbf{Y}}_{2}=m_{{\mathbf{Y}}_{1}}({\mathbf{X}},{\mathbf{W}}_{2})$ from model $M_{{\mathbf{Y}}_{1}}$ where ${\mathbf{W}}_{2}$ is another pseudo-random sequence. The ensemble model $M_{ensemble}$ is defined by ${\mathbf{Y}}_{2}=m_{ensemble}({\mathbf{X}},{\mathbf{W}}_{1},{\mathbf{W}}_{2})$.
The following observation is crucial to our main result: a good random number generator will almost always generate pseudo-random sequences ${\mathbf{W}}$ that belong to a typical set i.e. ${\mathbf{W}}$ that satisfies the simple AEP in Proposition 1.
Definition 4.
Typical String. An infinite text string ${\mathbf{Y}}$ (or more precisely a sequence of text strings ${\mathbf{Y}}_{N},~{}N=1,2,\dots$) is said to be typical for a model M if the log-perplexity of sufficiently long sub-strings converges to their empirical entropy for model M i.e. if $\lim_{N\rightarrow\infty}|l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})|=0$.
We illustrate these ideas for a simple Markov language model in Section A.2.
Proposition 3.
AEP for Perplexity. For a long text string ${\mathbf{Y}}$ i.e. $N\gg 1$ generated from a language model with a given prompt ${\mathbf{X}}$, with high probability, the log perplexity $l_{M}({\mathbf{Y}})$ is close to the empirical entropy $h_{M}({\mathbf{Y}})$. More precisely:
$\lim_{N\rightarrow\infty}\Pr[|l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (17)
Proof.
Let ${\mathbf{W}}$ denote the pseudo-random sequence that generated our string ${\mathbf{Y}}$ from model M. Thus we have ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$. With high probability, ${\mathbf{W}}$ is a typical sequence. Consider the ensemble model $M_{ensemble}$ for model M with pseudo-random sequences ${\mathbf{W}}_{1},~{}{\mathbf{W}}_{2}$ and let ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$. The classical AEP in Proposition 2 for model $M^{\prime}_{{\mathbf{Y}}_{1}}$ selected by the ensemble model requires that random strings ${\mathbf{Y}}^{\prime}_{N}$ generated by model $M^{\prime}_{{\mathbf{Y}}_{1}}$ must almost always satisfy
$\lim_{N\rightarrow\infty}|l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)|=0$ (18)
Since ${\mathbf{W}}$ is almost always a typical sequence, we can set ${\mathbf{W}}_{1}={\mathbf{W}}_{2}={\mathbf{W}}$, in which case the output of the ensemble model is exactly the same as the output ${\mathbf{Y}}$ of model M. Setting ${\mathbf{W}}_{1}\equiv{\mathbf{W}}$ we get ${\mathbf{Y}}_{1}\equiv{\mathbf{Y}}$ and
$h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)=h_{M^{\prime}_{{\mathbf{Y}}}}(N)\equiv h_{M}({\mathbf{Y}}_{N}).$ (19)
Finally setting ${\mathbf{W}}_{2}\equiv{\mathbf{W}}$ gives ${\mathbf{Y}}^{\prime}\equiv{\mathbf{Y}}$ and
$l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})=l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N}).$ (20)
Inserting (19, 20) into (18) proves the result. ∎
Proposition 3 asserts that with high probability a sufficiently long text string will be typical for the model that it was generated from. Note the crucial role in the above proof of the assumption that the pseudo-random sequence ${\mathbf{W}}$ is typical. We can take this construct further to establish some non-asymptotic results as well. We expect from (16) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ will be within a few standard deviations of its mean e.g. $h_{M}({\mathbf{Y}}_{N})\pm 3\lambda_{M}({\mathbf{Y}}_{N})$ even when $N$ is not large enough for asymptotics to kick in. Finally, we will show that the set of typical strings is vanishingly small compared to all grammatically correct strings.
Proposition 4.
Let $\mathcal{T}_{M}^{N}\subset\mathcal{T}^{N}$ denote the set of substrings ${\mathbf{Y}}_{N}$ of typical strings ${\mathbf{Y}}$ for a model M. The size of the “typical set” $\mathcal{T}_{M}^{N}$ is an exponentially vanishing fraction $02^{-\gamma N},~{}\gamma>$ 0 of all grammatically correct sequences.
Proof.
If we use top-k sampling, the model M generates $k$ possible grammatically correct values for each token $Y_{n}$. The number $T_{gram}$ of grammatically correct strings is then lower-bounded as $T_{gram}>k^{N}\equiv 2^{N\log_{2}k}$. Under the (mild) assumption that for a long string ${\mathbf{Y}}_{N}$, at least some fraction of the token distributions $p_{n}(y)$ are non-uniform, we have $h_{M}({\mathbf{Y}}_{N})\leq h_{max}<\log_{2}k$. By definition, the probability of model M generating string ${\mathbf{Y}}_{N}$ is $2^{-N~{}l_{M}({\mathbf{Y}}_{N})}$, so we have for the probability $P_{typ}$ of a typical string ${\mathbf{Y}}_{N}$ generated by model M:
$P_{typ}$ $\doteq 2^{-N~{}l_{M}({\mathbf{Y}}_{N})}\rightarrow 2^{-N~{}h_{M}({\mathbf{Y}}_{N})}>2^{-N~{}h_{max}}$ (21)
The total number of typical strings cannot exceed $\left|\mathcal{T}_{M}^{N}\right|\leq 2^{N~{}h_{max}}<T_{gram}~{}2^{-N\left(\log_{2}k-h_{max}\right)}$. ∎
Proposition 4 suggests that the typical sets of two different models, each set being vanishingly small, likely have zero overlap.
4 Experiments with Open-Source Language Model
We performed a series of experiments to verify the ideas described in Section 3. The basic idea is to use a local instance of the pre-trained GPT-2 model (Radford et al., 2019) to evaluate the empirical perplexity of generating a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions. Software code for the experiments is available at Bell and Mudumbai (2024).
4.1 Perplexity for GPT-2 of a String Generated by GPT-2
Our first set of experiments consisted of repeatedly generating a long text string from GPT-2 and then evaluating its log-perplexity, empirical entropy and log-deviations for the same GPT-2 model.
We initialized the prompt string and the seed of random number generator with fixed values to allow for later reproduction. We then used the top-k sampling with some fixed $k$ to generate the probability distribution for the next token. We normalized the probabilities of the top-k tokens to sum to unity. We then created and saved a table consisting of token ID, token string and selection probability of each of the top $k$ tokens. Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}=512$ total tokens which is the maximum number allowed by the GPT-2 model.
The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing. We ran this experiment many times for different seed values for the random number generator, different choices of $k$ for the top-k sampling and a number of different initial prompts.
For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3. Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
4.2 Perplexity for GPT-2 of an Arbitrary String
We also used a slightly modified version of the procedure described in Section 4.1 to calculate the log-perplexity and empirical entropy of an arbitrary string (that was not generated by GPT-2) on the GPT-2 model. We first choose a text string that we wish to analyze. We fix an initial fragment of the string as our prompt ${\mathbf{X}}$. We tokenize the remaining portion of the string excluding the prompt using the GPT-2 tokenizer. Let $Y_{n},~{}n=1\dots N_{max}$ denote the resulting sequence of tokens.
Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for GPT-2 using top-k sampling for the $n$’th token. We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 4.1. However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that GPT-2 will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^3] Note that our choice of top-k sampling is a compromise. The number of tokens in GPT-2 is quite large ($|\mathcal{T}|\approx 40,000$). If we exhaustively list the full probability distribution of all tokens, our experiment becomes very computationally expensive and slow.. When this happens, we replaced the lowest probability entry in the list of top-k tokens with the actual next token from our string along with its probability.
The normalized probability of each of the top-k tokens along with their token-ID and token strings, as well as the probability and ID of the actual next token are saved in a json file. This json file is then processed in exactly the same way as in Section 4.1 to compute the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the sub-string ${\mathbf{Y}}_{N}$ for the GPT-2 model for $N=1\dots N_{max}$.
4.3 Results and Discussion
Figure 2 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as a function of $N$ for three text strings generated by the same GPT-2 model using top-k sampling with $k=100$ for Fig. 2(a) and 2(b) and $k=4$ for Fig. 2(c). The plots also shows a range around the empirical entropy corresponding to ±$\pm\lambda_{N}({\mathbf{Y}}_{N})$ and ±$\pm 2\lambda_{N}({\mathbf{Y}}_{N})$. We see that in all three cases, the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently. We also see that the long-term average empirical entropy is around $3.5$ bits per token in Fig. 2(a) and Fig. 2(b), and $1.5$ bits per token in Fig. 2(c), which corresponds to approximately $2^{3.5}\approx 11$ alternative tokens to choose from for top-100 sampling and $2^{1.5}\approx 3$ tokens for top-4 sampling.
(a) GPT-2-generated string with top-100 sampling.
(b) GPT-2-generated string with top-100 sampling.
(c) GPT-2-generated string with top-4 sampling.
Figure 2: Log-perplexity and empirical entropy for GPT-2 of strings $\mathbf{Y}$ generated by the same model.
(a) Human written text.
(b) Claude 3 written text.
(c) Decreasing entropy.
Figure 3: Log-perplexity and empirical entropy for GPT-2 of selected text strings.
Figures 3(a) and 3(b) show the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as described in Section 4.2 for a string not generated by the GPT-2 model. The string in Fig. 3(a) is a paragraph written by the authors for an early draft of this manuscript, and the string in Fig. 3(b) was generated by the Claude 3 Opus model from Anthropic (2024). While our theory does not require non-convergence of $l_{M}({\mathbf{Y}}_{N})$ to $h_{M}({\mathbf{Y}}_{N})$, we do see from Fig. 3(a) and Fig. 3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ away from the empirical entropy $h_{M}({\mathbf{Y}}_{N})$. The contrast between Figs. 3 and 2 has obvious implications for AI text detection (Tang et al., 2024), model fingerprinting (Yu et al., 2021) and so on.
4.4 Anomalies and Limitations
While the results from our experiments described in Section 4 were always consistent with Proposition 3, they did reveal certain limitations of our setup that we now discuss.
For all three strings in Fig. 2, we see that $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$, but the log-deviations $\lambda_{M}({\mathbf{Y}}_{N})$ fail to become negligibly small even at $N=512$ suggesting that the GPT-2 model is not capable of generating strings sufficiently long for asymptotics to apply.
Fig. 3(c) shows an example of an output generated by GPT-2 for which the log-perplexity and empirical entropy both monotonically decrease with string length over many tokens. This suggests that the model repeatedly produces very low entropy token distributions $p_{n}(y)$, which happens for degenerate distributions where the probability mass is concentrated in a small number of tokens. In fact, on closer examination, we find that in these cases, the model is stuck in a state where the next token distribution just picks one token deterministically. This results in output text that gets stuck in a loop, repeating the same substring until $N_{max}$ has been reached.
Fundamentally, these limitations arise from cost and computational constraints. More powerful LLMs can generate longer texts and be less likely to get stuck in bad states that produce degenerate distributions and so on. In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti). Additional plots and more details about these experiments are presented in Section A.1.
5 Conclusions
We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property. While this paper focused narrowly on an exposition of the AEP property, it opens up many questions for further study. For instance, Proposition 4 suggests that different language models have distinct and unique “model signatures.” The contrast between Fig. 2 and Fig. 3 suggests potential applications for understanding and enhancing AI detection algorithms. As reported in Section 4.4, our experimental results are limited by computational constraints and can be strengthened by improved hardware and software. A broader question is to explore other statistical laws and constraints that generative models are bound by.
Appendix A Supplemental Material
A.1 Additional Experimental Context and Results
In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti). On this setup, each run of a text generation (i.e., running GPT-2 to generate $N=512$ tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds. We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.
Figure 4 shows additional results similar to Fig. 2 where we naively varied the seed value of the random number generator from 100-105. We present these to show that Fig. 2’s results are representative of the general case and were not cherry-picked. In all cases, $K=100$ and $N=512$. Further, should the community like to validate or expand upon this work, we have open sourced our repository [Bell and Mudumbai, 2024]. In addition to the source code, the supplemental material includes an /examples directory that includes the target text and the anomalous generated text for the plots in Fig. 3.
(a) Seed=100.
(b) Seed=101.
(c) Seed=102.
(d) Seed=103.
(e) Seed=104.
(f) Seed=105.
Figure 4: Log-perplexity and empirical entropy for GPT-2-generated string with top-100 sampling.
A.2 A Simple Markov Language Model
We now illustrate the ideas from Section 3 for a simple Markov language model, $M_{1}$ with a binary token space consisting of only two tokens $\mathcal{T}=\{t_{1},t_{2}\}$ and a parameter $\rho\in(0,1)$. The user prompt ${\mathbf{X}}$ consists of a single token which we can set to be $t_{1}$ without loss of generality: ${\mathbf{X}}=[t_{1}]$. The model $M_{1}$ generates a text string of length $N$ following a simple probabilistic algorithm: the $n$’th token $Y_{n}$ is set equal to the $n-1$’th token with probability $\rho$. The prompt serves as the initial token i.e. $Y_{0}\doteq t_{1}$. Let $h_{b}(\rho)\doteq-\rho\log_{2}\rho-(1-\rho)\log_{2}(1-\rho)$ denote the binary entropy function with argument $\rho$.
It is easy to see that $H(p_{n})\equiv h_{b}(\rho),~{}\forall n$ and $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(\rho)$ for any text string ${\mathbf{Y}}$. For $\rho=0.9$, $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(0.9)\approx 0.47$
Example 1.
Consider the string ${\mathbf{Y}}_{1}$ with all tokens $Y_{n}=t_{1},~{}\forall n=1\dots N$. The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(0.9)\approx 0.15$.
Example 2.
Consider the string ${\mathbf{Y}}_{2}$ with alternating token values: ${\mathbf{Y}}_{2}\equiv[t_{2},t_{1},t_{2},t_{1},t_{2},t_{1}\dots]$. The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(1-\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(0.1)\approx 3.3$.
Note that ${\mathbf{Y}}_{1}$ and ${\mathbf{Y}}_{2}$ are the most likely and least likely string of length $N$ respectively for $\rho=0.9$ (or indeed for any $\rho>0.5$). However, it is clear that neither string is typical for any reasonable $\epsilon$ with the perplexity of ${\mathbf{Y}}_{1}$ being abnormally low and that of ${\mathbf{Y}}_{2}$ abnormally high respectively.
Example 3.
Consider the string ${\mathbf{Y}}_{3}$ created by flipping a fair coin repeatedly to generate a vector whose elements are iid randomly selected with $\Pr(t_{1})=\Pr(t_{2})=0.5$. Since $Y_{3}$ is a random string, its perplexity $l_{M_{1}}({\mathbf{Y}}_{3})$ is also a random variable. However, for long strings i.e. $N\gg 1$, with high probability, $l_{M_{1}}({\mathbf{Y}}_{3})\approx-0.5\log_{2}\rho-0.5\log_{2}(1-\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{3})\approx 1.74$ which is incidentally the arithmetic mean of $l_{M_{1}}({\mathbf{Y}}_{1})$ and $l_{M_{1}}({\mathbf{Y}}_{2})$.
Clearly, with high probability, the string ${\mathbf{Y}}_{3}$ is not typical. Indeed, from these examples, we can see that typical strings for the model $M_{1}$ are precisely those that contain close to $(1-\rho)N$ token ’flips’ i.e. instances when $Y_{n+1}\neq Y_{n}$.

### v3

@@
 Abstract
-We propose a new asymptotic equipartition property for the perplexity of a large piece of text generated by a language model and present theoretical arguments for this property.
-Perplexity, defined as a inverse likelihood function, is widely used as a performance metric for training language models.
-Our main result states that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions.
-This means that language models are constrained to only produce outputs from a “typical set”, which we show, is a vanishingly small subset of all possible grammatically correct outputs.
-We present preliminary experimental results from an open-source language model to support our theoretical claims.
-This work has possible practical applications for understanding and improving “AI detection” tools and theoretical implications for the uniqueness, predictability and creative potential of generative models.
+We prove a new asymptotic equipartition property for the perplexity of long texts generated by a language model and present supporting experimental evidence from open-source models.
+Specifically we show that the logarithmic perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions.
+This defines a “typical set” that all long synthetic texts generated by a language model must belong to.
+We show that this typical set is a vanishingly small subset of all possible grammatically correct outputs.
+These results suggest possible applications to important practical problems such as (a) detecting synthetic AI-generated text, and (b) testing whether a text was used to train a language model.
+We make no simplifying assumptions (such as stationarity) about the statistics of language model outputs, and therefore our results are directly applicable to practical real-world models without any approximations.
+Language Models, Perplexity, Asymptotic Equipartition, Typical Sets
 1 Introduction and Motivation.
@@
 1.1 Contribution: Equipartition Property for Perplexity
-Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister and Cotterell, 2021).
-It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover and Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg and Keller, 2008; Cohen and Pakhomov, 2020) phenomena at least partially.
+Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister & Cotterell, 2021).
+It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover & Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg & Keller, 2008; Cohen & Pakhomov, 2020) phenomena at least partially.
 Many “AI detection” tools for identifying synthetic text from language models are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
-Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory.
+Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory.
 The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn.
-The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara and Morita, 2000; Timo et al., 2010).
+The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara & Morita, 2000; Timo et al., 2010).
 The outputs of non-trivial language models, however, cannot be reasonably modeled as a sequence of nearly-independent symbols (Brown et al., 2020).
+Our main contribution is to formulate and provide theoretical and experimental justification for a version of the AEP that applies to real-world practical LLMs without making simplifying assumptions about their statistics.
+A secondary contribution is to show possible applications of this theory to practical problems such as AI text detection and LLM dataset inference.
 2 Background and Definitions
@@
 We can write
-$P({\mathbf{Y}}_{N}|{\mathbf{X}})$ $=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
+$P$ $({\mathbf{Y}}_{N}|{\mathbf{X}})=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),$
+$~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
 Given a string ${\mathbf{Y}}$, open-source LLMs can be programmed to print out the distributions $p_{n}(Y_{n})$ from which its tokens were selected.
@@
 However, the conditional distributions $p_{n}(y)$ are not in general easily accessible, so while (2) is true for all generative models, it may only be useful for sequential models.
-The perplexity $p_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$.
+The perplexity $\mathrm{perp}_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$.
 It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
-$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(p_{M}({\mathbf{Y}}_{N})\right)\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
+$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(\mathrm{perp}_{M}({\mathbf{Y}}_{N})\right)$
+$\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
 2.2 A Toy Problem
@@
 $p_{\mathbf{A}}(y)$ $\doteq\frac{n_{\mathbf{A}}(y)}{N}\equiv\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}(A_{n}\equiv y),~{}\forall y\in\mathcal{T}$ (4)
-where $\mathbf{1}(.)$ denotes the indicator function and $n_{\mathbf{A}}(y)$ is the relative frequency of token $y$ in the (long) string ${\mathbf{A}}_{N}$.
+where $\mathbf{1}(.)$ denotes the indicator function and $n_{\mathbf{A}}(y)$ is the number of occurrences of token $y$ in the (long) string ${\mathbf{A}}_{N}$.
 The log-perplexity of string ${\mathbf{A}}_{N}$ for model A is:
-$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
+$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))$
+$\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
 where $H(\gamma_{1},\gamma_{2})\doteq-\sum_{y\in\mathcal{T}}\gamma_{1}(y)\log_{2}(\gamma_{2}(y))$ is the cross-entropy between two distributions $\gamma_{1}(y),~{}\gamma_{2}(y)$ over $y\in\mathcal{T}$.
@@
 Thus we have from (5):
-$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(\alpha)~{}{\mathrm{with~{}equality~{}iff~{}}}p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
-The simplest version of the classical Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory states that the log-perplexity of a long string ${\mathbf{A}}_{N}$ of iid symbols is almost always very close to the entropy $H(\alpha)$ of the distribution $\alpha(y)$ the symbols are drawn from.
-Proposition 1.
+$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(p_{\mathbf{A}})$
+${\mathrm{with~{}equality~{}iff~{}}}$ $p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
+The simplest version of the classical Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory states that the log-perplexity of a long string ${\mathbf{A}}_{N}$ of iid symbols is almost always very close to the entropy $H(\alpha)$ of the distribution $\alpha(y)$ the symbols are drawn from.
+Proposition 2.1.
 Simple AEP.
@@
 $\lim_{N\rightarrow\infty}\Pr[|l_{A}({\mathbf{A}}_{N})-H(\alpha)|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (7)
-Proposition 1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e.
+Proposition 2.1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e.
 convergence in probability.
 applied to the random variable $l_{M}({\mathbf{A}}_{N})$:
-$l_{M}({\mathbf{A}}_{N})$ $\rightarrow E\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
+$l_{M}({\mathbf{A}}_{N})$ $\rightarrow\operatorname*{{E}}\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
 From (5) and (6), we see that (8) itself is equivalent to $p_{\mathbf{A}}(y)\equiv\frac{n_{\mathbf{A}}(y)}{N}\rightarrow\alpha(y),~{}\forall y\in\mathcal{T}$ i.e.
@@
 Putting these observations together we have:
-$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(\alpha)\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
+$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(p_{\mathbf{A}})$
+$\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
 Intuitively, (9) states that a long string ${\mathbf{A}}$ generated from model A is likely to have a lower perplexity for model A than for any other model B.
 This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts.
-This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided that we accept the Shannon model of language i.e.
+This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided only that we accept the Shannon model of language i.e.
 the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961).
-2.3 A Naive Method for Detecting Text Generated by a Language Model
-Given a text string ${\mathbf{Y}}$, we seek to determine whether or not it was generated by a given model M.
-This is the problem of “white-box” detection of AI generated text where we assume knowledge of the specific language model that potentially generated the text.
-A naive method for doing this is as follows: compare the log-perplexity $l_{M}({\mathbf{Y}})$ of the text ${\mathbf{Y}}$ for model M with a threshold: if $l_{M}({\mathbf{Y}})<l_{th}$, we determine that ${\mathbf{Y}}$ is likely to have been generated by model M.
-Consider a long string ${\mathbf{B}}\doteq[B_{1},~{}B_{2},~{}\dots,~{}B_{N}]$ randomly generated from model B. Using the same argument as in Section 2.2, we have:
-$l_{B}({\mathbf{B}})$ $\equiv H(p_{\mathbf{B}},\beta)\rightarrow H(\beta)\leq H(p_{\mathbf{B}},\alpha)\equiv l_{A}({\mathbf{B}})$ (10)
-Note, however, that we have no basis to assert any relationship between $l_{A}({\mathbf{A}})$ and $l_{A}({\mathbf{B}})$.
-Indeed if $H(\beta)\ll H(\alpha)$ i.e.
-the distribution $\beta$ has significantly lower entropy than $\alpha$, it cannot be ruled out that $l_{A}({\mathbf{B}})<l_{A}({\mathbf{A}})$ i.e.
-a text-string generated by a different model has lower perplexity for a model than a text-string generated by itself.
-The mathematical properties of cross-entropy provide a rigorous basis for using perplexity to compare the compatibility of different models with a given text string.
-However, this theory does not support using perplexity to compare different text strings for compatibility with a given model.
+This same reasoning also underlies popular tools for LLM-related applications as we discuss in Section 5.
 2.3 A Modest Generalization
-This theoretical analysis relies on Proposition 1 which only applies to toy models that generate strings of iid tokens.
-Of course, the simple AEP in Proposition 1 has been extended in various ways in the literature such as the following version (see Tal et al.
-(2017), Appendix B.4), which we will make use of in the sequel.
-Proposition 2.
+This theoretical analysis relies on Proposition 2.1 which only applies to toy models that generate strings of iid tokens.
+Of course, the simple AEP in Proposition 2.1 has been extended in various ways in the literature such as the following version (see (Tal et al., 2017), Appendix B.4), which we will make use of in the sequel.
+Proposition 2.2.
 Generalized AEP.
@@
 $H(\alpha_{n}(y))<H_{max}<\infty,~{}\forall n$, the log perplexity $l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})$ of the string ${\mathbf{A}}^{\prime}_{N}$ as defined in (3) is close to the average entropy of the distributions $\alpha_{n}(y),~{}n=1,2,\dots,N$ with high probability:
-$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,~{}\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (11)
+$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,$
+$\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (10)
 The regularity condition that the entropies of $\alpha_{n}(y)$ are uniformly bounded is trivially true if the token dictionary $\mathcal{T}$ is a finite set: $H(p(y))\leq\log_{2}|\mathcal{T}|$ for any distribution $p(y)$ over $y\in\mathcal{T}$.
-While Proposition 2 does not lend itself to an intuitive interpretation in terms of the empirical distribution of the tokens $A^{\prime}_{n}$, it too is a direct consequence of the Law of Large Numbers applied to the log-perplexity random variable.
-Much of the analysis in Section 2.2 can be extended to generalized models like A’ in Proposition 2 that allows each token $A^{\prime}_{n}$ to be drawn from different distributions $\alpha_{n}(y)$.
+While Proposition 2.2 does not lend itself to an intuitive interpretation in terms of the empirical distribution of the tokens $A^{\prime}_{n}$, it too is a direct consequence of the Law of Large Numbers applied to the log-perplexity random variable.
+Much of the analysis in Section 2.2 can be extended to generalized models like $A^{\prime}$ in Proposition 2.2 that allows each token $A^{\prime}_{n}$ to be drawn from different distributions $\alpha_{n}(y)$.
 However, the tokens in these models must be drawn from fixed distributions independent of past tokens.
@@
 Specifically, given a model M and a string ${\mathbf{Y}}$, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{T}$.
-We will assume that the prompt ${\mathbf{X}}$ is fixed.
-Definition 1.
+We will assume that the prompt ${\mathbf{X}}$ is fixed and omit it from our notation in the sequel.
+Definition 3.1.
 The empirical entropy $h_{M}({\mathbf{Y}}_{N})$ of model M for string ${\mathbf{Y}}_{N}$ is defined as:
-$h_{M}({\mathbf{Y}}_{N})\doteq\frac{1}{N}\sum_{n=1}^{M}H(p_{n})$ (12)
+$h_{M}({\mathbf{Y}}_{N})\doteq\frac{1}{N}\sum_{n=1}^{N}H(p_{n})$ (11)
 where $H(p)\equiv-\sum_{y\in\mathcal{T}}p(y)\log_{2}(p(y))$ is the entropy of the distribution $p(y)$ over $y\in\mathcal{T}$.
-Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number.
+Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number.
 Indeed, we can show that
-$E\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv E\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$ (13)
+$\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv\operatorname*{{E}}\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$ (12)
 where the expectation is over random text strings ${\mathbf{Y}}_{N}$.
-Definition 2.
+Definition 3.2.
 Let the log-deviation $\lambda(p)$ of a probability distribution $p(y)$ over $y\in\mathcal{T}$ be defined as:
-$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}}~{}\mathrm{where}~{}S(p)\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (14)
+$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}},$
+$\mathrm{where}~{}S(p)$ $\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (13)
 Note that the entropy $H(p)$ and log-deviation $\lambda(p)$ are the mean and standard deviation of the log-likelihood random variable $l_{p}(y)\equiv-\log_{2}p(y)$ under distribution $p(y)$.
-Definition 3.
+Definition 3.3.
 The log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of a string ${\mathbf{Y}}_{N}$ for model $M$ is defined as:
@@
 Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string ${\mathbf{Y}}$ are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes.
-Preview of our main result.
-We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$.
-But this is very wrong - while the log-deviation $\lambda(p)$ of a fixed probability distribution $p(y)$ is a deterministic number, $\lambda_{M}({\mathbf{Y}}_{N})$ of a string ${\mathbf{Y}}_{N}$ is a random variable that takes different numerical values for different random text strings ${\mathbf{Y}}$.
-This considerably complicates our analysis.
-Fundamentally, their randomness arises from the fact that the distributions $p_{n}(y)$ are different for different strings ${\mathbf{Y}}$.
-This is because the tokens $Y_{n}$ in the string ${\mathbf{Y}}$ are not generated independently; instead future tokens strongly depend on earlier tokens.
-This is an essential property of modern language models that allows them to generate sophisticated text outputs.
-We now introduce the conceptual device of an auxiliary language model to get around this difficulty.
-3.1 Concept of an Auxiliary Language Model
-Let us consider a fixed string ${\mathbf{Y}}$.
-For this particular string ${\mathbf{Y}}$ and model M, we have a fixed set of distributions $p_{n}(y)$.
-Now define an auxiliary generative model $M^{\prime}_{\mathbf{Y}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~{}Y_{2}^{\prime},~{}\dots,~{}Y_{N}^{\prime},~{}\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $p_{n}(y)$.
-This construct of the auxiliary model $M^{\prime}_{\mathbf{Y}}$ for a fixed string ${\mathbf{Y}}$ allows us to work with fixed distributions $p_{n}(y)$ which are not themselves functions of a random string.
-In particular, note that the empirical entropy $h_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ and log-deviation $\lambda_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ for model $M^{\prime}_{\mathbf{Y}}$ are both deterministic functions of $N$ independent of the string ${\mathbf{Y}}^{\prime}_{N}$.
-Specifically, we have $h_{M^{\prime}_{\mathbf{Y}}}(N)\equiv h_{M}({\mathbf{Y}}_{N})$.
-Similarly, we have $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)\equiv\lambda_{M}({\mathbf{Y}}_{N})$ which, since we have fixed our special string ${\mathbf{Y}}$, is now also fixed.
-Indeed, it is easy to show that $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)$ is the standard-deviation of the random variable $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ the log-likelihood of the random string ${\mathbf{Y}}^{\prime}_{N}$.
-Then, by the Chebyshev Inequality (Cohen, 2015), we have:
-$\Pr\left(\left|l_{M}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{\mathbf{Y}}}(N)\right|>\epsilon\right)$ $\leq\frac{\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)}{\epsilon^{2}}$ (16)
-Clearly, for substrings ${\mathbf{Y}}_{N}$ of our special string ${\mathbf{Y}}$ that defines the auxiliary model $M^{\prime}_{\mathbf{Y}}$, we have by definition $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N})$.
-In addition, the entropy $H(P({\mathbf{Y}}^{\prime}_{N}))$ of the random strings ${\mathbf{Y}}^{\prime}_{N}$ output by auxiliary model $M^{\prime}_{\mathbf{Y}}$ is by definition $H(P({\mathbf{Y}}^{\prime}_{N}))\equiv h_{M}({\mathbf{Y}}_{N})$.
-The auxiliary model $M^{\prime}_{\mathbf{Y}}$ is much simpler than the general language model M.
-In particular, because the tokens $Y^{\prime}_{n}$ are generated independently of other tokens, the AEP in Proposition 2 applies to strings ${\mathbf{Y}}^{\prime}$ generated by the auxiliary model $M^{\prime}_{\mathbf{Y}}$.
-Indeed we note from (15) that the log-deviation $\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)$ vanishes for large $N$ under some regularity conditions on the distributions $p_{n}(y)$, in which case, the AEP for model $M^{\prime}_{\mathbf{Y}}$ follows from (16).
-3.2 Ensemble of Auxiliary Models
-We can think of a generative model M as a statistical ensemble of a large number of (simpler) auxiliary models.
-Each auxiliary model is capable of generating a random text string consisting of a sequence of independent tokens.
-We can then imagine choosing randomly from one of the auxiliary models.
-We now make this idea more precise.
-Consider an enumeration ${\mathbf{Y}}^{(1)},~{}{\mathbf{Y}}^{(2)},~{}\dots,~{}\dots,~{}{\mathbf{Y}}^{(N|\mathcal{T}|)}$ of the possible text strings ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$.
-Let $M_{{\mathbf{Y}}^{(n)}}$ denote the auxiliary model associated with model M and string ${\mathbf{Y}}^{(n)}$.
-A sequential generative model can be represented by a probability tree where each branch in the tree represents a new token being generated.
-The model M as well as each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can each be represented by such a probability tree.
-An example of a probability tree representing a generative model and a sample random string ${\mathbf{Y}}$ represented by a sequence of branches through the tree are shown in Fig.
-1 which also shows the probability tree for one possible auxiliary model $M^{\prime}_{\mathbf{Y}}$.
-Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens.
-Also shown in red is the probability tree of an auxiliary model for a string ${\mathbf{Y}}$ with $Y_{1}=3,~{}Y_{2}=1$ and so on.
-A generative model uses a pseudo-random number generator (PRNG) that produces a sequence of random symbols ${\mathbf{W}}$ to randomly choose a sequence of branches through its probability tree.
-As an example, suppose ${\mathbf{W}}=[W_{1},~{}W_{2},~{}\dots W_{N}~{}\dots]$ where the $W_{n}$ are independent and uniformly distributed in $[0,1]$ and a generative model uses $W_{n}$ to generate the n’th token $Y_{n}$ according to the distribution $p_{n}(y)$ using a sampling algorithm such as the classical inversion method Martino et al.
-(2018).
-Once we fix the PRNG and the sampling method, we can in principle write down a deterministic function to describe the generative model M as ${\mathbf{Y}}_{N}=m({\mathbf{X}},{\mathbf{W}}_{N})$ where recall that ${\mathbf{X}}$ is the fixed initial prompt.
-Each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can also similarly be described by deterministic functions ${\mathbf{Y}}_{N}=m_{{\mathbf{Y}}^{(n)}}({\mathbf{X}},{\mathbf{W}}_{N})$.
-We can then define a new generative model that we call the ensemble model $M_{ensemble}$ as follows.
-Pick a pseudo-random sequence ${\mathbf{W}}_{1}$ and generate a string ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$ from model M.
-Now consider the auxiliary model $M_{{\mathbf{Y}}_{1}}$ corresponding to string ${\mathbf{Y}}_{1}$.
-Generate a pseudo-random string ${\mathbf{Y}}_{2}=m_{{\mathbf{Y}}_{1}}({\mathbf{X}},{\mathbf{W}}_{2})$ from model $M_{{\mathbf{Y}}_{1}}$ where ${\mathbf{W}}_{2}$ is another pseudo-random sequence.
-The ensemble model $M_{ensemble}$ is defined by ${\mathbf{Y}}_{2}=m_{ensemble}({\mathbf{X}},{\mathbf{W}}_{1},{\mathbf{W}}_{2})$.
-The following observation is crucial to our main result: a good random number generator will almost always generate pseudo-random sequences ${\mathbf{W}}$ that belong to a typical set i.e.
-${\mathbf{W}}$ that satisfies the simple AEP in Proposition 1.
-Definition 4.
-Typical String.
-An infinite text string ${\mathbf{Y}}$ (or more precisely a sequence of text strings ${\mathbf{Y}}_{N},~{}N=1,2,\dots$) is said to be typical for a model M if the log-perplexity of sufficiently long sub-strings converges to their empirical entropy for model M i.e.
-if $\lim_{N\rightarrow\infty}|l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})|=0$.
-We illustrate these ideas for a simple Markov language model in Section A.2.
-Proposition 3.
+We are now ready to state our main result.
+Proposition 3.1.
 AEP for Perplexity.
@@
 $\lim_{N\rightarrow\infty}\Pr[|l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (15)
+We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$.
+This, however, is very wrong - for one thing $\lambda_{M}({\mathbf{Y}}_{N})$ is itself a random variable.
+In the sequel, we propose an Indifference Principle that allows us to carry through reasoning like the above to prove Proposition 3.1 and potentially other results.
+But first we present a direct proof of Proposition 3.1.
+3.1 Formal Proof: an LLN for LLMs
+We will use a version of the Weak Law of Large Numbers that applies to sequences of random variables with no restrictions on their dependence.
+Such a Law is presented in Appendix B (Proposition B.1).
+In fact, it turns out that a weaker version of Proposition B.1 namely Lemma B.2 is sufficient for our purposes.
+Proof of Proposition 3.1..
+Consider the sequence of conditional likelihood random variables $\{l_{n}\}$ defined as:
+$l_{n}$ $\doteq-\log_{2}\Pr\left(Y_{n}|{\mathbf{Y}}_{n-1}\right)\equiv-\log_{2}\left(p_{n}(Y_{n})\right)$ (16)
+We note that the random variables $Z_{n}\doteq l_{n}$ satisfy all of the conditions for Lemma B.2, if we impose the restriction that they are strictly bounded i.e.
+$|l_{n}|<L$, for some some finite $L$.
+This is equivalent to requiring that all non-zero token probabilities $p_{n}(y)$ are above some lower bound e.g.
+$p_{n}(y)\geq 10^{-10}$.
+In practice, this is easily satisfied by any practical LLM.
+Then from Lemma B.2, we have for any $\epsilon>0$:
+$\lim_{N\rightarrow\infty}\Pr\left(\left|\frac{1}{N}\sum_{n=1}^{N}\left(l_{n}-H(p_{n})\right)\right|>\epsilon\right)$ $=0$ (17)
+From the definition (16), we note that $\frac{1}{N}\sum_{n=1}^{N}l_{n}\equiv l_{M}({\mathbf{Y}}_{N})$, and likewise $\frac{1}{N}\sum_{n=1}^{N}H(p_{n})\equiv h_{M}({\mathbf{Y}}_{N})$.
+Combining these observations with (17) gives (15).
+∎
+3.2 Informal Proof: Roads not Traveled
+We now present an alternative proof of Proposition 3.1 using a more informal argument.
+Specifically, we introduce the conceptual device of an auxiliary language model that will allow us to apply the traditional Law of Large Numbers to a sequence of dependent random variables.
+Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens.
+Also shown in red is the probability tree of an auxiliary model $M^{\prime}_{\mathbf{S}}$ for a string ${\mathbf{S}}$ with $S_{1}=3,~{}S_{2}=1$ and so on.
+3.2.1 Concept of an Auxiliary Language Model
+Let us consider a fixed string ${\mathbf{S}}$.
+For this particular string ${\mathbf{S}}$ and model M, we have a fixed set of distributions $q_{n}(y)\doteq\Pr(S_{n}=y|{\mathbf{S}}_{n-1})$.
+Now define an auxiliary generative model $M^{\prime}_{\mathbf{S}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~{}Y_{2}^{\prime},~{}\dots,~{}Y_{N}^{\prime},~{}\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $q_{n}(y)$.
+If we represent a language model M as a probability tree, an auxiliary model $M^{\prime}_{\mathbf{S}}$ represents one specific subtree corresponding to the string ${\mathbf{S}}$ that defines the model.
+This is illustrated in Fig.
+1.
+The auxiliary model $M^{\prime}_{\mathbf{S}}$ is much simpler than the general language model M, because its output tokens $Y^{\prime}_{n}$ are generated independently of other tokens.
+Our informal proof of Proposition 3.1 is based on the following idea.
+Proposition 3.2.
+The Indifference Principle.
+For any random event $E$ defined entirely on the subtree of a particular string ${\mathbf{S}}$, $\Pr(E)$ will be the same for generative model M as for the auxiliary model $M^{\prime}_{\mathbf{S}}$.
+The idea behind Proposition 3.2 is that while calculating $\Pr(E)$ for some event $E$, we can change any part of the probability tree that do not affect the event $E$ (“the roads not traveled”) to make our calculation easier e.g.
+by allowing us to work with an auxiliary language model $M^{\prime}_{\mathbf{S}}$ which generates strings of independent tokens.
+For our purposes, Proposition 3.2 applies to all events in the $\sigma$-algebra generated by the conditional likelihood random variables $l_{n}$ as defined in (16).
+Informal Proof of Proposition 3.1..
+Using the Proposition 3.2, we can evaluate the distribution of random variables $l_{n}$ assuming the tokens $Y_{n}$ are generated by the auxiliary model $M^{\prime}_{\mathbf{S}}$ instead of the original model M.
+Then the random variables $l_{m},~{}l_{n}$ are independent of each other and $\operatorname*{{E}}\left[l_{n}\right]\equiv H\left(q_{n}\right)$.
+The weak Law of Large Numbers for the sequence of independent zero-mean random variables $l_{n}-H\left(q_{n}\right)$ immediately gives us (17).
+∎
+Proposition 3.2 allows us to use traditional statistical laws that apply to sequences of independent random variables directly without having to prove bespoke versions of those laws such as Proposition B.1.
+For instance, we may be able to use Proposition 3.2 to formulate a version of the Central Limit Theorem for $\frac{1}{\lambda_{M}({\mathbf{S}}_{N})}\left(l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right)$.
+We delegate a full development of these ideas to future work and limit ourselves to one simple non-asymptotic result to illustrate the method.
+3.3 Chebyshev Bound on Deviations
+We note that for the auxiliary model $M^{\prime}_{\mathbf{S}}$, the variance of the conditional log likelihood random variable $l_{n}$ is $\lambda^{2}(q_{n})$ as defined in (13).
+Assuming independence of $l_{n}$ using Proposition 3.2, we can show that the variance of $l_{M}({\mathbf{S}}_{N})$ is $\lambda^{2}_{M}({\mathbf{S}}_{N})$ as defined in (14).
+We can then write using the Chebyshev’s Inequality (Cohen, 2015) for any $\alpha>0$:
+$\Pr\left(\left|l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right|>\alpha\lambda_{M}({\mathbf{S}}_{N})\right)$ $\leq\frac{1}{\alpha^{2}}$ (18)
+3.4 Typical Sets are Vanishingly Small
+Let us now formally define the typical set $\mathcal{P}_{M}^{n}(\epsilon)\subset\mathcal{T}^{n}$ of typical strings ${\mathbf{Y}}_{n}$ for a model M with $\epsilon>0$:
+$\mathcal{P}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}$ (19)
+$\equiv\left\{{\mathbf{Y}}_{n}:\left|h_{M}({\mathbf{Y}}_{n})+\frac{1}{n}\log_{2}\left(\Pr({\mathbf{Y}}_{n})\right)\right|<\epsilon\right\}$
+Proposition 3.1 asserts that with high probability a stochastic model M is constrained to only output typical strings from this set $\mathcal{P}_{M}^{n}(\epsilon)$ for any $\epsilon>0$ if we consider sufficiently long strings.
+This observation has profound consequences.
+We will now show one such consequence, namely that the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ is very small:
+Proposition 3.3.
+The size of the typical set $\mathcal{P}_{M}^{n}$ is an exponentially vanishing fraction $02^{-\gamma n},~{}\gamma>$ 0 of all grammatically correct sequences.
+The argument below closely tracks standard information-theoretic results (Cover & Thomas, 2005a) about the size of the typical set. We start by restating Proposition B.1 informally as:
+$\Pr({\mathbf{Y}}_{n})$ $\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\rightarrow 2^{-nh_{M}({\mathbf{Y}}_{n})}$ (20)
+Let $T_{g}(n)$ be the total number of grammatically correct strings ${\mathbf{Y}}_{n}$ of length $n$.
+The average entropy $h_{M}({\mathbf{Y}})$ is a measure of the richness of the model that generated the string.
+Language models typically have a “temperature” parameter $T$ in the range $T\in[0,1]$ that can be used to tune the richness of the model’s output.
+Higher temperature means less “predictable” (i.e.
+less deterministic) and more “creative” (i.e.
+more random) outputs.
+This corresponds to higher values for $h_{M}({\mathbf{Y}})$.
+Consider the maximally rich model of length $n$ that has the highest possible entropy $h_{M}({\mathbf{Y}}_{n})\equiv h_{max}(n)$ which corresponds to choosing all possible (grammatically correct) strings of length $n$ with equal probability [^3] Note that this does not mean that the token distributions $p_{m}(y)$ are uniform over all valid tokens at each step $m$; indeed tokens corresponding to richer sub-trees as shown in Fig.
+1 should be chosen with higher probability to maximize the overall richness of the model’s output..
+By definition, for the maximally rich model, the probability of choosing any valid string is $2^{-nh_{max}(n)}$ and this means the number of possible strings of length $n$ is:
+$T_{g}(n)$ $\equiv 2^{nh_{max}(n)}$ (21)
+We can take (21) as the defining relation for $h_{max}(n)$.
+In practical LLMs, even at the highest temperature setting i.e.
+$T=1$, the token probability is typically concentrated in a small number of tokens and the token probabilities are determined by far more complex rules than the richness of their sub-trees.
+Let us formalize this as an assumption.
+Assumption 3.1.
+Practical LLMs are not entropy-maximizing. Any language model M has an entropy gap $\Delta h>0$ such that the following holds for all sufficiently long strings:
+$h_{M}({\mathbf{Y}}_{n})$ $<h_{max}(n)-\Delta h,~{}\forall{\mathbf{Y}}_{n}\in\mathcal{T}^{n}$ (22)
+Proof of Proposition 3.3.
+From (20), for $n\gg 1$, we have $\Pr({\mathbf{Y}}_{n})\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\to 2^{-nh_{M}({\mathbf{Y}}_{n})}>2^{-n(h_{max}(n)-\Delta h)}$ where we used Assumption 3.1 in the last step.
+Since this holds for any typical string of length $n$, we must have for small $\epsilon>0$ and correspondingly large $n$:
+$\left|\mathcal{P}_{M}^{n}(\epsilon)\right|$ $<2^{n\left(h_{max}(n)-\Delta h)\right)}\equiv 2^{-n\Delta h}\left|T_{g}(n)\right|$ (23)
+Thus, the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ represents an exponentially vanishing fraction of all grammatically correct strings. ∎
+We emphasize that all we needed to establish Proposition 3.3 was Assumption 3.1 which only requires that the model assigns higher probability to some sentences than to others.
+3.5 Over-Typical Sets are Vanishingly Smaller Still
+The obverse of Proposition 3.3 is that almost all grammatically correct strings are non-typical for a given language model M.
+A string can be non-typical in two different ways depending on whether its perplexity is abnormally high (“under-typical”) or abnormally low (“over-typical”).
+Formally, we define the set $\mathcal{V}_{M}^{n}(\epsilon)$ of over-typical strings of length $n$ for a model M as:
+$\mathcal{V}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left(h_{M}({\mathbf{Y}}_{n})-l_{M}({\mathbf{Y}}_{n})\right)>\epsilon\right\}$
+$\equiv\left\{{\mathbf{Y}}_{n}:\Pr({\mathbf{Y}}_{n})>2^{-n\left(h_{M}({\mathbf{Y}}_{n})-\epsilon\right)}\right\}$ (24)
+Proposition 3.4.
+The size of the over-typical set $\mathcal{V}_{M}^{n}(\epsilon)$ is an exponentially vanishing fraction of all grammatically correct sequences.
 Proof.
-Let ${\mathbf{W}}$ denote the pseudo-random sequence that generated our string ${\mathbf{Y}}$ from model M.
-Thus we have ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$.
-With high probability, ${\mathbf{W}}$ is a typical sequence.
-Consider the ensemble model $M_{ensemble}$ for model M with pseudo-random sequences ${\mathbf{W}}_{1},~{}{\mathbf{W}}_{2}$ and let ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$.
-The classical AEP in Proposition 2 for model $M^{\prime}_{{\mathbf{Y}}_{1}}$ selected by the ensemble model requires that random strings ${\mathbf{Y}}^{\prime}_{N}$ generated by model $M^{\prime}_{{\mathbf{Y}}_{1}}$ must almost always satisfy
-$\lim_{N\rightarrow\infty}|l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)|=0$ (18)
-Since ${\mathbf{W}}$ is almost always a typical sequence, we can set ${\mathbf{W}}_{1}={\mathbf{W}}_{2}={\mathbf{W}}$, in which case the output of the ensemble model is exactly the same as the output ${\mathbf{Y}}$ of model M.
-Setting ${\mathbf{W}}_{1}\equiv{\mathbf{W}}$ we get ${\mathbf{Y}}_{1}\equiv{\mathbf{Y}}$ and
-$h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)=h_{M^{\prime}_{{\mathbf{Y}}}}(N)\equiv h_{M}({\mathbf{Y}}_{N}).$ (19)
-Finally setting ${\mathbf{W}}_{2}\equiv{\mathbf{W}}$ gives ${\mathbf{Y}}^{\prime}\equiv{\mathbf{Y}}$ and
-$l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})=l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N}).$ (20)
-Inserting (19, 20) into (18) proves the result. ∎
-Proposition 3 asserts that with high probability a sufficiently long text string will be typical for the model that it was generated from.
-Note the crucial role in the above proof of the assumption that the pseudo-random sequence ${\mathbf{W}}$ is typical.
-We can take this construct further to establish some non-asymptotic results as well.
-We expect from (16) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ will be within a few standard deviations of its mean e.g.
-$h_{M}({\mathbf{Y}}_{N})\pm 3\lambda_{M}({\mathbf{Y}}_{N})$ even when $N$ is not large enough for asymptotics to kick in.
-Finally, we will show that the set of typical strings is vanishingly small compared to all grammatically correct strings.
-Proposition 4.
-Let $\mathcal{T}_{M}^{N}\subset\mathcal{T}^{N}$ denote the set of substrings ${\mathbf{Y}}_{N}$ of typical strings ${\mathbf{Y}}$ for a model M.
-The size of the “typical set” $\mathcal{T}_{M}^{N}$ is an exponentially vanishing fraction $02^{-\gamma N},~{}\gamma>$ 0 of all grammatically correct sequences.
-Proof.
-If we use top-k sampling, the model M generates $k$ possible grammatically correct values for each token $Y_{n}$.
-The number $T_{gram}$ of grammatically correct strings is then lower-bounded as $T_{gram}>k^{N}\equiv 2^{N\log_{2}k}$.
-Under the (mild) assumption that for a long string ${\mathbf{Y}}_{N}$, at least some fraction of the token distributions $p_{n}(y)$ are non-uniform, we have $h_{M}({\mathbf{Y}}_{N})\leq h_{max}<\log_{2}k$.
-By definition, the probability of model M generating string ${\mathbf{Y}}_{N}$ is $2^{-N~{}l_{M}({\mathbf{Y}}_{N})}$, so we have for the probability $P_{typ}$ of a typical string ${\mathbf{Y}}_{N}$ generated by model M:
-$P_{typ}$ $\doteq 2^{-N~{}l_{M}({\mathbf{Y}}_{N})}\rightarrow 2^{-N~{}h_{M}({\mathbf{Y}}_{N})}>2^{-N~{}h_{max}}$ (21)
-The total number of typical strings cannot exceed $\left|\mathcal{T}_{M}^{N}\right|\leq 2^{N~{}h_{max}}<T_{gram}~{}2^{-N\left(\log_{2}k-h_{max}\right)}$. ∎
-Proposition 4 suggests that the typical sets of two different models, each set being vanishingly small, likely have zero overlap.
-4 Experiments with Open-Source Language Model
+For any ${\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)$, we have from (24) and (22):
+$\Pr({\mathbf{Y}}_{n})$ $>2^{-nh_{max}(n)}\times 2^{n(\Delta h+\epsilon)}$ (25)
+Therefore, we must have
+$\left|\mathcal{V}_{M}^{n}(\epsilon)\right|$ $<\left|T_{g}(n)\right|2^{-n(\Delta h+\epsilon)}$
+where we used the definition of $h_{max}$ in (21). ∎
+We have shown that almost all long grammatically correct strings are under-typical for any language model M. We illustrate these ideas for a simple Markov language model in Appendix A.
+4 Experiments with Open-Source Models
 We performed a series of experiments to verify the ideas described in Section 3.
-The basic idea is to use a local instance of the pre-trained GPT-2 model (Radford et al., 2019) to evaluate the empirical perplexity of generating a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions.
-Software code for the experiments is available at Bell and Mudumbai (2024).
-4.1 Perplexity for GPT-2 of a String Generated by GPT-2
-Our first set of experiments consisted of repeatedly generating a long text string from GPT-2 and then evaluating its log-perplexity, empirical entropy and log-deviations for the same GPT-2 model.
+The basic idea is to use a local instance of a pre-trained open-source model to evaluate the empirical entropy of a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions.
+We ran experiments on the smaller GPT-2 model (Radford et al., 2019) as well as the more recent, larger and more capable Llama 3.1 8B (Touvron et al., 2023) model.
+4.1 Perplexity of Self-Generated Strings
+Our first set of experiments consisted of repeatedly generating a long text string from a model (GPT-2 or Llama 3.1 8B) and then evaluating its log-perplexity, empirical entropy and log-deviations for the same model.
 We initialized the prompt string and the seed of random number generator with fixed values to allow for later reproduction.
-We then used the top-k sampling with some fixed $k$ to generate the probability distribution for the next token.
+We then used top-k sampling with $k=100$ to generate the probability distribution for the next token.
 We normalized the probabilities of the top-k tokens to sum to unity.
 We then created and saved a table consisting of token ID, token string and selection probability of each of the top $k$ tokens.
-Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}=512$ total tokens which is the maximum number allowed by the GPT-2 model.
+Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}$ total tokens.
+We want the number of tokens $N_{max}$ to be large enough to observe the asymptotic behavior of the perplexity of the generated string.
+In practice, $N_{max}$ is constrained by certain limitations of the models as we discuss in the sequel.
 The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing.
-We ran this experiment many times for different seed values for the random number generator, different choices of $k$ for the top-k sampling and a number of different initial prompts.
-For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3.
+We ran this experiment many times for different seed values for the random number generator, different choices of sampling distributions and a number of different initial prompts.
+For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3.1.
 Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
-4.2 Perplexity for GPT-2 of an Arbitrary String
-We also used a slightly modified version of the procedure described in Section 4.1 to calculate the log-perplexity and empirical entropy of an arbitrary string (that was not generated by GPT-2) on the GPT-2 model.
+4.2 Perplexity of Externally Generated Strings
+We also used a slightly modified version of the procedure described in Section 4.1 to calculate the log-perplexity and empirical entropy of an arbitrary string (that was not generated by the model itself) on the GPT-2 and Llama 3.1 8B models.
 We first choose a text string that we wish to analyze.
 We fix an initial fragment of the string as our prompt ${\mathbf{X}}$.
-We tokenize the remaining portion of the string excluding the prompt using the GPT-2 tokenizer.
+We tokenize the remaining portion of the string excluding the prompt using the model’s tokenizer.
 Let $Y_{n},~{}n=1\dots N_{max}$ denote the resulting sequence of tokens.
-Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for GPT-2 using top-k sampling for the $n$’th token.
+Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for the model using top-k with $k=100$ sampling for the $n$’th token.
 We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 4.1.
-However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that GPT-2 will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^3] Note that our choice of top-k sampling is a compromise.
-The number of tokens in GPT-2 is quite large ($|\mathcal{T}|\approx 40,000$).
+However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that our model will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^4] Note that our choice of top-k sampling is a compromise.
+The number of tokens in GPT-2 or Llama is quite large ($|\mathcal{T}|\approx 40,000$).
 If we exhaustively list the full probability distribution of all tokens, our experiment becomes very computationally expensive and slow..
@@
 4.3 Results and Discussion
-Figure 2 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as a function of $N$ for three text strings generated by the same GPT-2 model using top-k sampling with $k=100$ for Fig.
-2(a) and 2(b) and $k=4$ for Fig.
-2(c).
-The plots also shows a range around the empirical entropy corresponding to ±$\pm\lambda_{N}({\mathbf{Y}}_{N})$ and ±$\pm 2\lambda_{N}({\mathbf{Y}}_{N})$.
-We see that in all three cases, the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently.
-We also see that the long-term average empirical entropy is around $3.5$ bits per token in Fig.
-2(a) and Fig.
-2(b), and $1.5$ bits per token in Fig.
-2(c), which corresponds to approximately $2^{3.5}\approx 11$ alternative tokens to choose from for top-100 sampling and $2^{1.5}\approx 3$ tokens for top-4 sampling.
-(a) GPT-2-generated string with top-100 sampling.
-(b) GPT-2-generated string with top-100 sampling.
-(c) GPT-2-generated string with top-4 sampling.
-Figure 2: Log-perplexity and empirical entropy for GPT-2 of strings $\mathbf{Y}$ generated by the same model.
-(a) Human written text.
-(b) Claude 3 written text.
-(c) Decreasing entropy.
-Figure 3: Log-perplexity and empirical entropy for GPT-2 of selected text strings.
-Figures 3(a) and 3(b) show the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as described in Section 4.2 for a string not generated by the GPT-2 model.
-The string in Fig.
-3(a) is a paragraph written by the authors for an early draft of this manuscript, and the string in Fig.
-3(b) was generated by the Claude 3 Opus model from Anthropic (2024).
-While our theory does not require non-convergence of $l_{M}({\mathbf{Y}}_{N})$ to $h_{M}({\mathbf{Y}}_{N})$, we do see from Fig.
+Figure 2 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 or Llama 3.1 8B model as a function of $N$ for several text strings generated by the same model.
+The plots also shows the range $h_{M}({\mathbf{Y}}_{N})\pm\lambda_{M}({\mathbf{Y}}_{N})$ and $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$.
+We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently in all cases.
+The asymptotic convergence is more clearly seen in the long strings from the Llama 3.1 8B model [^5] The GPT-2 model cannot generate similarly long texts.
+in Fig.
+2(c), 2(d) .
+We find that very long synthetic texts such as those in Fig.
+2(c), 2(d) can be ungrammatical, repetitive and generally of poor quality, as reported elsewhere (Holtzman et al., 2020).
+However, Proposition 3.1 holds regardless of the quality of the text as seen in Fig.
+2.
+(a) GPT-2, $N_{max}=512$.
+(b) Llama 3.1 8B, $N_{max}=1024$.
+(c) Llama 3.1 8B, $N_{max}=10240$.
+(d) Llama 3.1 8B, $N_{max}=10240$.
+Figure 2: Log-perplexity and empirical entropy for GPT-2 or Llama 3.1 8B of self-generated text, top-100 sampling.
+Figure 3 shows the log-perplexity and empirical entropy for the Llama 3.1 8B model for two long, non-Llama generated strings.
+The text used in Fig.
+3(a) is an excerpt from a recent article (Pelly, 2025) in Harper’s Magazine, a respected literary publication.
+The string used in Fig.
+3(b) was generated by ChatGPT o1 Pro (OpenAI, 2024).
+(a) Human writing.
+(b) Text from ChatGPT o1 pro.
+Figure 3: Log-perplexity and empirical entropy for Llama 3.1 8B of non-Llama-generated strings $\mathbf{Y}$.
+We see from Fig.
 3(a) and Fig.
-3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ away from the empirical entropy $h_{M}({\mathbf{Y}}_{N})$.
-The contrast between Figs.
-3 and 2 has obvious implications for AI text detection (Tang et al., 2024), model fingerprinting (Yu et al., 2021) and so on.
-4.4 Anomalies and Limitations
-While the results from our experiments described in Section 4 were always consistent with Proposition 3, they did reveal certain limitations of our setup that we now discuss.
-For all three strings in Fig.
-2, we see that $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$, but the log-deviations $\lambda_{M}({\mathbf{Y}}_{N})$ fail to become negligibly small even at $N=512$ suggesting that the GPT-2 model is not capable of generating strings sufficiently long for asymptotics to apply.
-Fig.
-3(c) shows an example of an output generated by GPT-2 for which the log-perplexity and empirical entropy both monotonically decrease with string length over many tokens.
-This suggests that the model repeatedly produces very low entropy token distributions $p_{n}(y)$, which happens for degenerate distributions where the probability mass is concentrated in a small number of tokens.
-In fact, on closer examination, we find that in these cases, the model is stuck in a state where the next token distribution just picks one token deterministically.
-This results in output text that gets stuck in a loop, repeating the same substring until $N_{max}$ has been reached.
-Fundamentally, these limitations arise from cost and computational constraints.
-More powerful LLMs can generate longer texts and be less likely to get stuck in bad states that produce degenerate distributions and so on.
-In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti).
-Additional plots and more details about these experiments are presented in Section A.1.
+3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ for Llama 3.1 8B is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ larger than the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ showing that the corresponding strings are clearly under-typical for the Llama 3.1 8B model.
+While this non-convergence is not required by our theory - we have not considered whether the converse to Proposition 3.1 is true - it is strongly suggested by Propositions 3.3 and 3.4 if we assume that the strings in Fig.
+3 are randomly sampled from the set of all grammatically correct texts.
+4.4 More Catholic than the Pope
+(a) Alice in Wonderland excerpt.
+(b) The Great Gatsby excerpt.
+Figure 4: Log-perplexity and empirical entropy for Llama 3.1 8B of excerpts from literary classics.
+This brings us to the remarkable plots in Fig.
+4 which shows the log-perplexity of excerpts from classic texts for Llama 3.1 8B, specifically from Alice in Wonderland and The Great Gatsby in Figs.
+4(a) and 4(b) respectively.
+The average entropy for the token distributions in Fig.
+4(a) is an astonishingly small $0.11$ bits, which means that the model on average assigns a very high probability ($\sim 99\%$) to the precise sequence of tokens from the excerpt for its predicted next tokens.
+In other words, this excerpt has been memorized by the Llama 3.1 8B model.
+The same observation also applies to Fig.
+4(b).
+We also see from both plots that the log-perplexity is several standard deviations below the empirical entropy making the excerpts slightly, but unambiguously, over-typical for the Llama-8B model: the model judges these texts to be more perfect than its own creations!
+From Proposition 3.4, we can conclude that these excerpts are among the small number of very special strings to be so memorized by the model.
+Clearly, these texts were part of the training set of the Llama-8B model and they were esteemed highly by its cost function.
+Interestingly, the over-typicality of these excerpts means that it is pretty much impossible for the Llama-8B model itself to generate them verbatim - even with a $99\%$ probability that each next token will exactly match the excerpt, a randomly generated string can be expected to diverge from the excerpt after only about $100$ tokens.
+Additional plots and more details about these experiments are presented in Appendix C.1.
+5 Practical Applications
+Consider the problem of AI text detection i.e.
+determining whether a string ${\mathbf{Y}}$ was generated by a model M.
+The intuition behind (9) suggests a simple method for doing this: check if the perplexity is small enough.
+Specifically, declare ‘yes’ if $l_{M}({\mathbf{Y}})<l_{th}$ for a threshold $l_{th}$.
+The same test can also be used for the problem of LLM dataset inference i.e.
+determining whether a string ${\mathbf{Y}}$ was used as part of the training dataset for the model M.
+While, this idea is not new and indeed variations of it are behind popular solutions for both problems (Mitchell et al., 2023; Maini et al., 2024), our results in this paper provide a rigorous theoretical foundation for it.
+It also provides a principled way to choose the classification threshold $l_{th}$ which previously had to be computed empirically (Carlini et al., 2021).
+As an example, with $l_{th}=h_{M}({\mathbf{Y}})+3\lambda_{M}({\mathbf{Y}})$, Fig.
+2 shows that we can accurately distinguish between synthetic text generated by Llama 3.1 8B and non-Llama text such as in Fig.
+3 even with relatively short strings of $1000$ tokens.
+We can also provide performance guarantees.
+Indeed, a simple lower-bound on the probability of a false negative (i.e.
+failing to detect a Llama-generated string) can be obtained from (18).
+Similar comments apply to the dataset inference problem: while it is well-known that text strings memorized by an LLM have low perplexity, our theory allows us to rigorously derive threshold values and performance guarantees.
+Extrapolating from Fig.
+4, we posit that if a long excerpt from a human-written book is over-typical for a language model, it is a near-statistical certainty that the book was part of the training data set for that model.
 6 Conclusions
 We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property.
-While this paper focused narrowly on an exposition of the AEP property, it opens up many questions for further study.
-For instance, Proposition 4 suggests that different language models have distinct and unique “model signatures.” The contrast between Fig.
-2 and Fig.
-3 suggests potential applications for understanding and enhancing AI detection algorithms.
-As reported in Section 4.4, our experimental results are limited by computational constraints and can be strengthened by improved hardware and software.
-A broader question is to explore other statistical laws and constraints that generative models are bound by.
-Appendix A Supplemental Material
-A.1 Additional Experimental Context and Results
-In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti).
-On this setup, each run of a text generation (i.e., running GPT-2 to generate $N=512$ tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds.
-We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.
-Figure 4 shows additional results similar to Fig.
-2 where we naively varied the seed value of the random number generator from 100-105.
-We present these to show that Fig.
-2’s results are representative of the general case and were not cherry-picked.
-In all cases, $K=100$ and $N=512$.
-Further, should the community like to validate or expand upon this work, we have open sourced our repository [Bell and Mudumbai, 2024].
-In addition to the source code, the supplemental material includes an /examples directory that includes the target text and the anomalous generated text for the plots in Fig.
-3.
-(a) Seed=100.
-(b) Seed=101.
-(c) Seed=102.
-(d) Seed=103.
-(e) Seed=104.
-(f) Seed=105.
-Figure 4: Log-perplexity and empirical entropy for GPT-2-generated string with top-100 sampling.
-A.2 A Simple Markov Language Model
+While this paper focused on a rigorous exposition of the AEP property, it opens up many questions for further study.
+An important open problem is to design practical algorithms for AI text detection and dataset inference using the typical set concept.
+Tighter non-asymptotic bounds on deviations that improve on (18) will significantly benefit practical applications by providing better performance guarantees.
+More fundamentally, Proposition 3.3 leads naturally to the question of how much do the typical sets of different models overlap.
+The theory presented in this paper is neutral on this question which must be answered empirically and both possible answers have interesting implications.
+No overlap between typical sets means that each model can be associated with a unique statistical signature.
+On the other hand, independently trained models converging on a common set of typical strings, would be a notable linguistic phenomenon, though it is hard to speculate on what this might mean.
+The AEP is also just one example of the broader observation that generative AI models are stochastic machines subject to statistical laws and the discovery of other statistical regularities in the outputs of generative models is an important topic for future work.
+Appendix A A Simple Markov Language Model
 We now illustrate the ideas from Section 3 for a simple Markov language model, $M_{1}$ with a binary token space consisting of only two tokens $\mathcal{T}=\{t_{1},t_{2}\}$ and a parameter $\rho\in(0,1)$.
@@
 For $\rho=0.9$, $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(0.9)\approx 0.47$
-Example 1.
+Example A.1.
 Consider the string ${\mathbf{Y}}_{1}$ with all tokens $Y_{n}=t_{1},~{}\forall n=1\dots N$.
@@
 For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(0.9)\approx 0.15$.
-Example 2.
+Example A.2.
 Consider the string ${\mathbf{Y}}_{2}$ with alternating token values: ${\mathbf{Y}}_{2}\equiv[t_{2},t_{1},t_{2},t_{1},t_{2},t_{1}\dots]$.
@@
 However, it is clear that neither string is typical for any reasonable $\epsilon$ with the perplexity of ${\mathbf{Y}}_{1}$ being abnormally low and that of ${\mathbf{Y}}_{2}$ abnormally high respectively.
-Example 3.
+Example A.3.
 Consider the string ${\mathbf{Y}}_{3}$ created by flipping a fair coin repeatedly to generate a vector whose elements are iid randomly selected with $\Pr(t_{1})=\Pr(t_{2})=0.5$.
@@
 instances when $Y_{n+1}\neq Y_{n}$.
+Appendix B A Generalized Law of Large Numbers for Random Variables with Arbitrary Dependence
+Consider a infinitely long string ${\mathbf{Y}}=[Y_{1}~{}Y_{2}~{}\dots~{}Y_{n}~{}\dots]$ consisting of a sequence of symbols $Y_{n}\in\mathcal{Y}$ generated stochastically from a finite dictionary $\mathcal{Y}$.
+We will denote the substring consisting of the first $N$ symbols as ${\mathbf{Y}}_{N}\equiv[Y_{1}~{}Y_{2}~{}\dots Y_{N}]$ and the joint distribution of the symbols $P({\mathbf{Y}}_{N})$.
+We make no assumptions about the distribution $P({\mathbf{Y}}_{N})$ from which the symbols $Y_{n}$ are generated.
+In particular, we do not assume that the symbols are independent or weakly dependent even asymptotically.
+We also do not assume stationarity i.e.
+the marginal distributions of symbols $Y_{n}$ and $Y_{m}$ can be completely different.
+We also do not assume any Markovian or conditional independence properties.
+B.1 Conditional Means and Sample Means for Sequences of Dependent Random Variables
+Consider a sequence of random variables ${\mathbf{Z}}=[Z_{1}~{}Z_{2}~{}\dots~{}Z_{n}~{}\dots]$ whose $n$’th element $Z_{n}$ is a function of the first $n$ symbols $Y_{1},Y_{2},\dots,Y_{n}$ i.e.
+$Z_{n}\doteq g_{n}({\mathbf{Y}}_{n})$ and $g_{n}:\mathcal{Y}^{n}\rightarrow\mathbb{R}$ are functions that we will assume to be uniformly bounded i.e.
+$|g_{n}({\mathbf{Y}}_{n})|\leq G<\infty,~{}\forall n\in\mathbb{Z}^{+},{\mathbf{Y}}_{n}\in\mathcal{Y}^{n}$.
+We define the conditional mean and conditional variance of the random variables $Z_{n}$:
+$\mu_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)g_{n}({\mathbf{Y}}_{n})\equiv\operatorname*{{E}}\left[Z_{n}|{\mathbf{Y}}_{n-1}\right]$ (26)
+$V_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)\left(g_{n}({\mathbf{Y}}_{n})-\mu_{n}^{({\mathbf{Y}})}\right)^{2}\equiv\operatorname*{{E}}\left[Z_{n}^{2}|{\mathbf{Y}}_{n-1}\right]-\left(\mu_{n}^{({\mathbf{Y}})}\right)^{2}$ (27)
+where the distribution $p^{({\mathbf{Y}})}_{n}(y)\doteq\Pr(Y_{n}=y|{\mathbf{Y}}_{n-1})$.
+Since the random variables $Z_{n}$ are discrete-valued and bounded, they have bounded moments for any distribution:
+$|\mu_{n}^{({\mathbf{Y}})}|\leq G,~{}$ $V_{n}^{({\mathbf{Y}})}\leq G^{2}$ (28)
+A more formal statement.
+Let $\mathcal{F}_{n}$ be the $\sigma$-algebra defined by ${\mathbf{Y}}_{n}\equiv[Y_{1}~{}Y_{2}~{}\dots~{}Y_{n}]$ i.e.
+$\mathcal{F}_{n}$ is a collection that includes all (mathematically reasonable) events defined on the first $n$ symbols.
+Then the random variables $Z_{n}$ as we have defined them are defined by the condition that $Z_{m}$ is $\mathcal{F}_{m}$-measurable, and $\mu_{m}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{m}\big{|}\mathcal{F}_{m-1}\right]$.
+We need some intermediate results before presenting our main result. First, we define the partial sums of the sequence $Z_{n}$ and conditional means:
+$S_{n}$ $\doteq\sum_{m=1}^{n}Z_{m},~{}\mu_{S_{n}}^{({\mathbf{Y}})}\doteq\sum_{m=1}^{n}\mu_{m}^{({\mathbf{Y}})}$ (29)
+Clearly, $\operatorname*{{E}}\left[Z_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{n}^{({\mathbf{Y}})}\right]$ and $\operatorname*{{E}}\left[S_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{S_{n}}^{({\mathbf{Y}})}\right]$.
+Lemma B.1.
+The variance of the zero-mean random variable $S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}$ can be written as sums of contributions from each of the $Z_{m}$’s. Specifically, we have:
+$V_{S_{n}}$ $\doteq\operatorname*{{E}}\left[\left(S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right)^{2}\right]=\sum_{m=1}^{n}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]$ (30)
+Proof.
+We will use the method of induction to prove (30). For $n=1$, (30) holds trivially. Now, let’s assume (30) holds for $n=l$ i.e. assume that:
+$V_{S_{l}}$ $\equiv\sum_{m=1}^{l}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]$ (31)
+Now consider $S_{l+1}\equiv S_{l}+Z_{l+1}$. Its variance can be written as:
+$V_{S_{l+1}}$ $=\operatorname*{{E}}\left[\left(S_{l}+Z_{l+1}-\mu_{S_{l}}^{({\mathbf{Y}})}-\mu_{l+1}^{({\mathbf{Y}})}\right)^{2}\right]$ (32)
+$=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]-2\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]$ (33)
+We will now show that the last term in (33) is zero. We apply the law of iterated expectations to write:
+$\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]=\operatorname*{{E}}_{{\mathbf{Y}}_{l}}\left[\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\big{|}{\mathbf{Y}}_{l}\right]\right]$ (34)
+$=\operatorname*{{E}}_{{\mathbf{Y}}_{l}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\operatorname*{{E}}\left[\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\big{|}{\mathbf{Y}}_{l}\right]\right]\equiv 0$ (35)
+where we used the fact that $S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}$ is a deterministic function of ${\mathbf{Y}}_{l}$ and $\mu^{({\mathbf{Y}})}_{l+1}\equiv\operatorname*{{E}}\left[Z_{l+1}|{\mathbf{Y}}_{l}\right]$.
+Remark.
+An alternative way to arrive at (35) is to treat $\mu_{l+1}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{l+1}\big{|}{\mathbf{Y}}_{l}\right]$ as an estimate of $Z_{l+1}$ given ${\mathbf{Y}}_{l}$, and then use the famous Orthogonality Principle to argue that the estimation error $\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)$ must be uncorrelated with any function of ${\mathbf{Y}}_{l}$.
+More formally, $\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)$ is $\mathcal{F}_{l}$-measurable and is therefore orthogonal to $\left(Z_{l+1}-\operatorname*{{E}}\left[Z_{l+1}|\mathcal{F}_{l}\right]\right)$.
+Using (35), we can now simplify (33) to:
+$V_{S_{l+1}}$ $=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]\equiv\sum_{m=1}^{l+1}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]$ (36)
+where we used the induction assumption (31). We have shown that if (30) holds for $n=l$, then (30) must also hold for $n=l+1$, thus completing the induction. ∎
+Lemma B.2.
+The sequence $\frac{1}{n}\left(S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right)$ converges to zero in probability:
+$\lim_{n\rightarrow\infty}\Pr\left(\frac{1}{n}\left|S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~{}\forall\epsilon>0$ (37)
+Proof.
+Using Lemma B.1, we can write:
+$V_{S_{n}}$ $=\sum_{m=1}^{n}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]\leq nG^{2}$ (38)
+using the simple bounds in (28).
+The variance of $\frac{1}{n}\left(S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right)$ is simply $\frac{1}{n^{2}}V_{S_{n}}\leq\frac{G^{2}}{n}\rightarrow 0$ as $n\rightarrow\infty$.
+Equation (37) now follows immediately using the Chebyshev Inequality (Cohen, 2015).
+∎
+Lemma B.2 states that the sample means of the two random sequences $\{Z_{n}\}$ and $\{\mu^{({\mathbf{Y}})}_{n}\}$ must be close to each other asymptotically.
+Since we have made very few assumptions about the random variables involved, we cannot assume that either of the sequences $\{Z_{n}\}$ or $\{\mu^{({\mathbf{Y}})}_{n}\}$ themselves converge in any sense to an asymptotic limit.
+Neither can we assume that their sample means separately converge to a limit.
+With an additional assumption stating the existence of one such limit, we can establish a Weak Law of Large Numbers generalized to its most extreme i.e.
+a statement that the sample mean of a sequence of random variables $\{Z_{n}\}$ with arbitrary dependence converges asymptotically to a suitably defined average quantity in probability.
+Proposition B.1.
+A Weak Law of Large Numbers.
+If the average of the conditional means converges to a limit i.e.
+$\mu^{({\mathbf{Y}})}\equiv\lim\limits_{n\rightarrow\infty}\frac{1}{n}\sum\limits_{m=1}^{n}\mu_{m}^{({\mathbf{Y}})}$ exists, then the sequence $\frac{1}{n}S_{n}$ converges in probability to $\mu^{({\mathbf{Y}})}$:
+$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=1}^{n}Z_{n}$ $\to_{p}\mu^{({\mathbf{Y}})}~{}~{}\mathrm{or}~{}\lim_{n\to\infty}\Pr\left(\left|\frac{1}{n}S_{n}-\mu^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~{}\forall\epsilon>0$ (39)
+Proof.
+Combining Lemma B.2 with the observation $\frac{1}{n}\mu_{S_{n}}^{({\mathbf{Y}})}\rightarrow\mu^{({\mathbf{Y}})}$ completes the proof. ∎
+B.2 Discussion
+Proposition B.1 is remarkable because unlike traditional versions of the LLN, it makes no assumptions whatsoever about the dependence or correlation between the symbols $Y_{n}$.
+It does not even require for instance that the correlation between $Y_{1}$ and $Y_{n}$ should decrease as $n$ becomes large.
+The two extreme cases of strong dependence and complete independence for the symbols $Y_{n}$ turn out to be two simple special cases for our proposed LLN:
+  • If the symbols $Y_{n}$ are independent, Proposition B.1 simply reduces to the classical weak LLN.
+  • If the symbols are perfectly correlated, the $Z_{m}$ become deterministic functions of the first token, and the LLN then becomes a trivial observation.
+Appendix C Supplemental Material
+C.1 Additional Experimental Context and Results
+In this work, all experiments were ran on a high-performance computing server running Linux (Ubuntu).
+AMD EPYC 7413 24-Core Processor, 48 logical cores, 96 GB RAM, four NVIDIA A100-SXM4-80GB GPUs.
+On this setup, each run of a text generation (i.e., running GPT-2 to generate Nmax = 512 tokens, Llama-8B to generate Nmax = 1024 tokens and Nmax = 10240 tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds to minutes, with each experiment taking about 90 minutes to run in the case of 10,240 steps.
+We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.3.
+Figure 5 shows additional results similar to Fig.
+2 where we naively varied the seed value of the random number generator from $209-214$ for Llama-8B.
+We present these to show that Fig.
+2’s results are representative of the general case and were not cherry-picked.
+In all cases, $K=100$.
+(a) Seed=209.
+(b) Seed=210.
+(c) Seed=211.
+(d) Seed=212.
+(e) Seed=213.
+(f) Seed=214.
+Figure 5: Log-perplexity and empirical entropy for Llama 3.1 of strings generated by the same model with top-100 sampling.

### v4

@@
 Abstract
-We prove a new asymptotic equipartition property for the perplexity of long texts generated by a language model and present supporting experimental evidence from open-source models.
+We prove a new asymptotic unequipartition property for the perplexity of long texts generated by a language model and present supporting experimental evidence from open-source models.
 Specifically we show that the logarithmic perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions.
 This defines a “typical set” that all long synthetic texts generated by a language model must belong to.
-We show that this typical set is a vanishingly small subset of all possible grammatically correct outputs.
-These results suggest possible applications to important practical problems such as (a) detecting synthetic AI-generated text, and (b) testing whether a text was used to train a language model.
+We refine the concept of ”typical set” to include only grammatically correct texts.
+We then show that this refined typical set is a vanishingly small subset of all possible grammatically correct texts for a very general definition of grammar.
+This means that language models are strongly constrained in the range of their possible behaviors and outputs.
 We make no simplifying assumptions (such as stationarity) about the statistics of language model outputs, and therefore our results are directly applicable to practical real-world models without any approximations.
-Language Models, Perplexity, Asymptotic Equipartition, Typical Sets
-1 Introduction and Motivation.
+We discuss possible applications of the typical set concept to problems such as detecting synthetic texts and membership inference in training datasets.
+1 Introduction and Background.
 Consider a generative model, defined as an algorithm that takes a user input $\mathbf{X}$} and produces an output $\mathbf{Y}$} that statistically resembles data from a natural source.
@@
 State-of-the-art LLMs (Anthropic, 2024; OpenAI, 2024) are now able to produce detailed and information-rich text outputs such as entire screenplays and book-length manuscripts from a short and simple user prompt.
-Furthermore, LLMs produce text that can imitate human language well-enough to pass the Turing test (Pinar Saygin et al., 2000) i.e.
-resembles text created by humans well enough to be convincing to human observers.
+They can also pass the Turing test (Pinar Saygin et al., 2000) i.e.
+imitate human-written text well enough to be convincing to human observers.
 In this work, we argue that an LLM is strongly constrained by statistical laws.
 Specifically, we show that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions.
-This means that any language model is constrained to only output text strings from a typical set, which is an exponentially vanishing subset of all possible grammatically correct strings.
-1.1 Contribution: Equipartition Property for Perplexity
-Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister & Cotterell, 2021).
-It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover & Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg & Keller, 2008; Cohen & Pakhomov, 2020) phenomena at least partially.
-Many “AI detection” tools for identifying synthetic text from language models are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
-Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory.
+This means that any language model is constrained to only output text strings from a typical set, which we show, is an exponentially vanishing subset of all possible grammatically correct strings.
+This involves a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory and refining it for language models to account for the concept of grammatical correctness.
+This work touches on ideas such as entropy, equipartition theorems and computational linguistics/natural language processing (CL/NLP), each of which boast of a vast body of previous work.
+We now present a brief literature survey to place our work in context.
+1.1 Equipartition Theorems
 The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn.
-The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara & Morita, 2000; Timo et al., 2010).
-The outputs of non-trivial language models, however, cannot be reasonably modeled as a sequence of nearly-independent symbols (Brown et al., 2020).
-Our main contribution is to formulate and provide theoretical and experimental justification for a version of the AEP that applies to real-world practical LLMs without making simplifying assumptions about their statistics.
-A secondary contribution is to show possible applications of this theory to practical problems such as AI text detection and LLM dataset inference.
-2 Background and Definitions
-Let M be a generative model described by ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$, where the output ${\mathbf{Y}}$ consisting of a (potentially infinite) string of tokens ${\mathbf{Y}}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}~{}\dots]$ is a deterministic function of user prompt ${\mathbf{X}}$ and a pseudo-random sequence ${\mathbf{W}}$.
-Each token is chosen from a finite set of tokens $Y_{n}\in\mathcal{T}$.
-The model M can produce a number of different output strings ${\mathbf{Y}}$ for the same user prompt ${\mathbf{X}}$ corresponding to different values of the pseudo-random sequence ${\mathbf{W}}$.
-This defines a probability distribution $P({\mathbf{Y}}|{\mathbf{X}})$, or more formally, a sequence of probability distributions $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ where ${\mathbf{Y}}_{N}\doteq[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ is the substring of ${\mathbf{Y}}$ consisting of the first $N$ tokens.
-2.1 Sequential LLM
+The AEP itself can be thought of as a consequence of the Law of Large Numbers (LLN) (Idele, 2018) that asserts the convergence of a log-likelihood random variable to a limiting “entropy rate” (Algoet & Cover, 1988).
+A weak AEP for iid sequences with a finite alphabet was originally introduced by Shannon (1948) followed by strong versions for stationary, ergodic sequences by McMillan (1953) and Breiman (1957), extensions to infinite alphabets by Chung (1961), and continuous-valued sequences in Barron (1985).
+Thus, the AEP is also commonly referred to as the Shannon-McMillan-Breiman (SMB) Theorem Hamdan et al.
+(2008).
+Generalizations of the SMB Theorem (Moy, 1961) in the form of ergodic theorems are now available for many classes of stationary sequences (Pollicott & Yuri, 1998).
+Extending the AEP for iid sequences to a non-stationary sequence of independent, but non-identically distributed symbols is straightforward (Tang et al., 2024).
+AEPs for word-valued functions of stationary sequences are considered in (Nishiara & Morita, 2000; Timo et al., 2010).
+However, beyond this, the literature on the AEP for non-stationary sequences is much more limited.
+Closest to our work is the analysis in Yang & Liu (2004) and Wang & Yang (2016) that consider a very general class of non-homogeneous Markov processes, but assume the existence of a unique limiting distribution for the Markov process to establish a constant entropy rate.
+Our key point of departure is to not require equal partitions: without stationarity or similarly restrictive assumptions on the statistics of the process, the likelihoods of long sequences do not all become asymptotically equal.
+Crucially, we show that the the typical sets defined by the resulting Un-Equipartition Property still retain certain essential properties that make the typical set concept so powerful for stationary processes.
+1.2 Computational Models of Language
+Statistical analysis of languages in its modern form dates back to the work of (Shannon, 1951) on the entropy of written English.
+These statistical ideas competed (Chomsky, 1959) with structural theories of language (Chomsky, 2014) in the early days of modern linguistics.
+Early methods such as n-gram models (Manning & Schütze, 1999) evolved into more mathematically sophisticated methods such as Hidden Markov Models (HMMs) (Rabiner, 1989).
+The statistical approach has now become dominant after the recent introduction of neural network architectures (Bengio et al., 2003), culminating in the development of powerful large language models (LLMs) based on transformer networks (Vaswani et al., 2017).
+While we know that LLMs can generate complex and coherent text outputs, their representation of language structures and grammar remains a topic of research Reizinger et al.
+(2024); Mészáros et al.
+(2024).
+Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for training language models (Meister & Cotterell, 2021).
+It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and entropy (Cover & Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg & Keller, 2008; Cohen & Pakhomov, 2020) phenomena at least partially.
+Many “AI detection” tools are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
+1.3 Contribution: An Un-Equipartition Property for Unnatural Language
+Our main contributions are (a) to formulate and provide theoretical and experimental justification for a version of the AEP that applies to real-world practical LLMs without any approximations, and (b) use this AEP to define a concept of typical set specifically adapted for language models.
+While our focus is on the theoretical development, we also briefly discuss possible applications to practical problems such as AI text detection and LLM dataset inference.
+Unlike computational linguistics, we do not study natural language; instead our exclusive focus is on unnatural language i.e.
+synthetic text created by LMMs.
+Natural languages are shaped by complex social, psychological and biological processes Pierce (1968) and they can only ever be approximated by computational models Bar-Hillel (1962).
+In contrast, LLMs are machines whose internals, while complex, are in theory completely knowable.
+Thus, it is possible for us to derive exact laws that govern the outputs of LLMs in a way that is not possible to do for natural language.
+However, this requires us to be disciplined about not making any assumptions or approximations about the statistics of LLM outputs.
+1.4 Notation
+A sequence of symbols is denoted by boldface, uppercase identifiers e.g.
+${\mathbf{A}},~{\mathbf{B}}$.
+These can be fixed, constant sequences or random sequences.
+The vector consisting of the first $N$ elements of a sequence $\mathbf{A}$} is denoted by ${\mathbf{A}}_{N}\equiv[A_{1},~A_{2},~\dots,~A_{N}]$.
+All random variables in this work take discrete-values from a finite set.
+The probability of an event $E$ is $\Pr(E)$, the distribution or probability mass function of a random variable $A_{m}$ is $p(A_{m})$.
+Sets are denoted by calligraphic symbols e.g.
+$\mathcal{A}$.
+The binary entropy of a distribution $p(A_{m})$ is $H(p(A_{m}))$ [^1] We will always speak of the entropy of distributions and avoid speaking of the entropy of random variables..
+The symbol $\equiv$ denotes a simple identity while == denotes a mathematical inference and $\doteq$ a definition.
+2 Definitions and Problem Statement
+Let M be a generative model whose output $\mathbf{Y}$} consisting of a sequence of tokens ${\mathbf{Y}}=[Y_{1},~Y_{2},~\dots Y_{N}~\dots]$, chosen from a finite token dictionary $Y_{n}\in\mathcal{Y}$, is a stochastic function of user prompt $\mathbf{X}$}.
+The model M is defined by a probability distribution $p({\mathbf{Y}}|{\mathbf{X}})$, or more formally, a sequence of probability distributions $p({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{Y}^{N}$ where ${\mathbf{Y}}_{N}\doteq[Y_{1},~Y_{2},~\dots Y_{N}]$ is the substring of $\mathbf{Y}$} consisting of the first $N$ tokens.
 Practical implementations of LLMs specify the probability distribution iteratively (Radford et al., 2018):
@@
 We can write
-$P$ $({\mathbf{Y}}_{N}|{\mathbf{X}})=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),$
-$~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
+$p$ $({\mathbf{Y}}_{N}|{\mathbf{X}})=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),~\mathrm{where}~p_{n}(Y_{n})=p(Y_{n}|Y_{1},~Y_{2},~Y_{n-1},{\mathbf{X}})$ (2)
 Given a string $\mathbf{Y}$}, open-source LLMs can be programmed to print out the distributions $p_{n}(Y_{n})$ from which its tokens were selected.
 Specifically, given a user prompt $\mathbf{X}$} and a string of tokens ${\mathbf{Y}}_{N}\equiv[Y_{1},~Y_{2},~\dots,~Y_{N}]$, it is possible to get a complete listing of the distributions $p_{n}(Y_{n}),~n=1\dots N$.
-Note that complete knowledge of the $p_{n}(y)$ is not the same as complete knowledge of $P({\mathbf{Y}}|{\mathbf{X}})$ even for a fixed user prompt ${\mathbf{X}}$.
-As an example, the former requires knowledge of $p_{3}(y)\equiv P(y|Y_{1},Y_{2},{\mathbf{X}})$ for different values of $y\in\mathcal{T}$, but only conditioned on the specific values of $Y_{1},~{}Y_{2}$ contained in one specific string ${\mathbf{Y}}$.
 Remark.
@@
 However, the conditional distributions $p_{n}(y)$ are not in general easily accessible, so while (2) is true for all generative models, it may only be useful for sequential models.
-The perplexity $\mathrm{perp}_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$.
+The perplexity $\mathrm{perp}_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a text string ${\mathbf{Y}}_{N}=[Y_{1},~Y_{2},~\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string $\mathbf{Y}$}.
 It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
-$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(\mathrm{perp}_{M}({\mathbf{Y}}_{N})\right)$
-$\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
+$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(\mathrm{perp}_{M}({\mathbf{Y}}_{N})\right)\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
 2.1 A Toy Problem
-Let $\alpha(y),~{}\beta(y)$ be two fixed probability distributions over the (discrete) set $\mathcal{T}$ of tokens.
+Let $\alpha(y),~\beta(y)$ be two fixed probability distributions over the (discrete) set $\mathcal{Y}$ of tokens.
 Consider a toy problem involving two language models A and B, that each generate a string of tokens ${\mathbf{Y}}=[Y_{1},~Y_{2},~\dots]$ where each token is generated iid from the distribution $\alpha(y)$ and $\beta(y)$ respectively.
@@
 Consider a long string ${\mathbf{A}}_{N}\doteq[A_{1},~A_{2},~\dots,~A_{N}],~N\gg 1$ randomly generated from model A.
-Let $p_{\mathbf{A}}(y)$ denote the empirical distribution of $y\in\mathcal{T}$ in the string ${\mathbf{A}}_{N}$:
-$p_{\mathbf{A}}(y)$ $\doteq\frac{n_{\mathbf{A}}(y)}{N}\equiv\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}(A_{n}\equiv y),~{}\forall y\in\mathcal{T}$ (4)
+Let $p_{\mathbf{A}}(y)$ denote the empirical distribution of $y\in\mathcal{Y}$ in the string ${\mathbf{A}}_{N}$:
+$p_{\mathbf{A}}(y)$ $\doteq\frac{n_{\mathbf{A}}(y)}{N}\equiv\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}(A_{n}\equiv y),~\forall y\in\mathcal{Y}$ (4)
 where $\mathbf{1}(.)$ denotes the indicator function and $n_{\mathbf{A}}(y)$ is the number of occurrences of token $y$ in the (long) string ${\mathbf{A}}_{N}$.
 The log-perplexity of string ${\mathbf{A}}_{N}$ for model A is:
-$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))$
-$\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
-where $H(\gamma_{1},\gamma_{2})\doteq-\sum_{y\in\mathcal{T}}\gamma_{1}(y)\log_{2}(\gamma_{2}(y))$ is the cross-entropy between two distributions $\gamma_{1}(y),~{}\gamma_{2}(y)$ over $y\in\mathcal{T}$.
-It is well-known $H(\gamma_{1},\gamma_{2})\geq H(\gamma_{1})$ with equality when $\gamma_{1}\equiv\gamma_{2}$, where $H(\gamma)\doteq-\sum_{y\in\mathcal{T}}\gamma(y)\log_{2}(\gamma(y))$ is the entropy of distribution $\gamma$.
+$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))\equiv-\sum_{y\in\mathcal{Y}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
+where $H(\gamma_{1},\gamma_{2})\doteq-\sum_{y\in\mathcal{Y}}\gamma_{1}(y)\log_{2}(\gamma_{2}(y))$ is the cross-entropy between distributions $\gamma_{1}(y),~\gamma_{2}(y)$ over $y\in\mathcal{Y}$.
+It is well-known $H(\gamma_{1},\gamma_{2})\geq H(\gamma_{1})$ with equality when $\gamma_{1}\equiv\gamma_{2}$, where $H(\gamma)\doteq-\sum_{y\in\mathcal{Y}}\gamma(y)\log_{2}(\gamma(y))$ is the entropy of distribution $\gamma$.
 Thus we have from (5):
-$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(p_{\mathbf{A}})$
-${\mathrm{with~{}equality~{}iff~{}}}$ $p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
+$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(p_{\mathbf{A}}){\mathrm{with~equality~iff~}}p_{\mathbf{A}}(y)\equiv\alpha(y),~\forall y\in\mathcal{Y}$ (6)
 The simplest version of the classical Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory states that the log-perplexity of a long string ${\mathbf{A}}_{N}$ of iid symbols is almost always very close to the entropy $H(\alpha)$ of the distribution $\alpha(y)$ the symbols are drawn from.
 Proposition 2.1.
-Simple AEP.
-For a long text string ${\mathbf{A}}_{N}=[A_{1},~{}A_{2},~{}\dots,~{}A_{N}]$ of iid tokens $A_{n}$ drawn from a distribution $\alpha(y)$, the log perplexity $l_{A}({\mathbf{A}}_{N})$ as defined in (3) is close to the entropy of the distribution $H(\alpha)$ of the distribution $\alpha(y)$ with high probability:
-$\lim_{N\rightarrow\infty}\Pr[|l_{A}({\mathbf{A}}_{N})-H(\alpha)|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (7)
-Proposition 2.1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e.
-convergence in probability.
-applied to the random variable $l_{M}({\mathbf{A}}_{N})$:
-$l_{M}({\mathbf{A}}_{N})$ $\rightarrow\operatorname*{{E}}\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
-From (5) and (6), we see that (8) itself is equivalent to $p_{\mathbf{A}}(y)\equiv\frac{n_{\mathbf{A}}(y)}{N}\rightarrow\alpha(y),~{}\forall y\in\mathcal{T}$ i.e.
-the empirical distribution $p_{\mathbf{A}}(y)$ for long strings ${\mathbf{A}}$ is very close to $\alpha(y)$ with high probability.
+A Simple Cross-AEP.
+For a long text string ${\mathbf{A}}_{N}=[A_{1},~A_{2},~\dots,~A_{N}]$ of iid tokens $A_{n}$ drawn from a distribution $\alpha(y)$, the log perplexity $l_{B}({\mathbf{A}}_{N})$ for a model B as defined in (3) almost always approaches the cross-entropy of the distributions $H(\alpha,\beta)$:
+$\lim_{N\to\infty}l_{B}({\mathbf{A}}_{N})$ $\equiv H(\alpha,\beta),~\mathrm{and}~\lim_{N\to\infty}l_{A}({\mathbf{A}}_{N})\equiv H(\alpha)$ (7)
+From (5) and (6), we see that (7) itself is equivalent to $p_{\mathbf{A}}(y)\equiv\frac{n_{\mathbf{A}}(y)}{N}\rightarrow\alpha(y),~\forall y\in\mathcal{Y}$ i.e.
+the empirical distribution $p_{\mathbf{A}}(y)$ for long strings $\mathbf{A}$} almost always converges to $\alpha(y)$.
 Putting these observations together we have:
-$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(p_{\mathbf{A}})$
-$\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
-Intuitively, (9) states that a long string ${\mathbf{A}}$ generated from model A is likely to have a lower perplexity for model A than for any other model B.
+$\lim_{N\rightarrow\infty}l_{A}({\mathbf{A}}_{N})$ $\equiv H(\alpha)\leq H(\alpha,\beta)\equiv\lim_{N\rightarrow\infty}l_{B}({\mathbf{A}}_{N})$ (8)
+Intuitively, (8) states that a long string ${\mathbf{A}}_{N}$ generated from model A is likely to have a lower perplexity for model A than for any other model B.
 This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts.
@@
 the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961).
-This same reasoning also underlies popular tools for LLM-related applications as we discuss in Section 5.
+Testing for abnormally high or low perplexity is an important tool for LLM applications as we discuss in Section 6.
 2.2 A Modest Generalization
-This theoretical analysis relies on Proposition 2.1 which only applies to toy models that generate strings of iid tokens.
-Of course, the simple AEP in Proposition 2.1 has been extended in various ways in the literature such as the following version (see (Tal et al., 2017), Appendix B.4), which we will make use of in the sequel.
+The simple AEP in Proposition 2.1 has been extended in various ways in the literature such as the following version (see Tal et al. (2017), Appendix B.4), which we will make use of in the sequel.
 Proposition 2.2.
@@
 $H(\alpha_{n}(y))<H_{max}<\infty,~\forall n$, the log perplexity $l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})$ of the string ${\mathbf{A}}^{\prime}_{N}$ as defined in (3) is close to the average entropy of the distributions $\alpha_{n}(y),~n=1,2,\dots,N$ with high probability:
-$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,$
-$\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (10)
-The regularity condition that the entropies of $\alpha_{n}(y)$ are uniformly bounded is trivially true if the token dictionary $\mathcal{T}$ is a finite set: $H(p(y))\leq\log_{2}|\mathcal{T}|$ for any distribution $p(y)$ over $y\in\mathcal{T}$.
+$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-h_{A^{\prime}}(N)|>\epsilon]\equiv 0,~\forall\epsilon>0,\mathrm{where}~h_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (9)
+The regularity condition that the entropies of $\alpha_{n}(y)$ are uniformly bounded is trivially true if the token dictionary $\mathcal{Y}$ is a finite set: $H(p(y))\leq\log_{2}|\mathcal{Y}|$ for any distribution $p(y)$ over $y\in\mathcal{Y}$.
 While Proposition 2.2 does not lend itself to an intuitive interpretation in terms of the empirical distribution of the tokens $A^{\prime}_{n}$, it too is a direct consequence of the Law of Large Numbers applied to the log-perplexity random variable.
@@
 This is still quite trivial compared to modern language models where the output tokens depend in highly complex and sophisticated ways on past tokens.
-3 An Equipartition Property for Perplexity
+3 An Un-Equipartition Property for Perplexity
 We now propose a generalization of the theory described in Section 2.1 to a more interesting class of models.
 Consider a language model M and a random infinitely long text string $\mathbf{Y}$} generated by M, whose probabilities for a given prompt $\mathbf{X}$} is described by (2).
-Specifically, given a model M and a string ${\mathbf{Y}}$, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{T}$.
+Specifically, given a model M and a string $\mathbf{Y}$}, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{Y}$.
 We will assume that the prompt $\mathbf{X}$} is fixed and omit it from our notation in the sequel.
@@
 $h_{M}({\mathbf{Y}}_{N})\doteq\frac{1}{N}\sum_{n=1}^{N}H(p_{n})$ (10)
-where $H(p)\equiv-\sum_{y\in\mathcal{T}}p(y)\log_{2}(p(y))$ is the entropy of the distribution $p(y)$ over $y\in\mathcal{T}$.
-Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number.
-Indeed, we can show that
-$\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv\operatorname*{{E}}\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$ (12)
-where the expectation is over random text strings ${\mathbf{Y}}_{N}$.
+where $H(p)\equiv-\sum_{y\in\mathcal{Y}}p(y)\log_{2}(p(y))$ is the entropy of the distribution $p(y)$ over $y\in\mathcal{Y}$.
+Note that we make no assumptions about the existence of an “entropy rate” $\lim_{N\rightarrow\infty}h_{M}({\mathbf{Y}}_{N})$ for the model $M$. However, we can show that
+$\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv\operatorname*{{E}}\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(\Pr({\mathbf{Y}}_{N})\right)$ (11)
+where $H\left(\Pr({\mathbf{Y}}_{N})\right)$ is the entropy of the random text strings ${\mathbf{Y}}_{N}$.
 Definition 3.2.
-Let the log-deviation $\lambda(p)$ of a probability distribution $p(y)$ over $y\in\mathcal{T}$ be defined as:
-$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}},$
-$\mathrm{where}~{}S(p)$ $\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (13)
+Let the log-deviation $\lambda(p)$ of a probability distribution $p(y)$ over $y\in\mathcal{Y}$ be defined as:
+$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}},~\mathrm{where}~S(p)\doteq\sum_{y\in\mathcal{Y}}p(y)\left(\log_{2}p(y)\right)^{2}$ (12)
 Note that the entropy $H(p)$ and log-deviation $\lambda(p)$ are the mean and standard deviation of the log-likelihood random variable $l_{p}(y)\equiv-\log_{2}p(y)$ under distribution $p(y)$.
@@
 $\lambda_{M}({\mathbf{Y}}_{N})$ $\doteq\frac{1}{N}\sqrt{\sum_{n=1}^{N}\lambda^{2}(p_{n})}$ (13)
-We can again interpret $H(p_{n}),~{}\lambda(p_{n})$ as the conditional mean and standard deviation of $-\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$.
+We can again interpret $H(p_{n}),~\lambda(p_{n})$ as the conditional mean and standard deviation of $\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$.
 Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string $\mathbf{Y}$} are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes.
@@
 Proposition 3.1.
-AEP for Perplexity.
+Un-Equipartition Property for Perplexity.
 For a long text string $\mathbf{Y}$} i.e.
@@
 This, however, is very wrong - for one thing $\lambda_{M}({\mathbf{Y}}_{N})$ is itself a random variable.
-In the sequel, we propose an Indifference Principle that allows us to carry through reasoning like the above to prove Proposition 3.1 and potentially other results.
+In the sequel, we show an informal argument to carry through reasoning like the above.
 But first we present a direct proof of Proposition 3.1.
@@
 We will use a version of the Weak Law of Large Numbers that applies to sequences of random variables with no restrictions on their dependence.
-Such a Law is presented in Appendix B (Proposition B.1).
-In fact, it turns out that a weaker version of Proposition B.1 namely Lemma B.2 is sufficient for our purposes.
+Such a Law can be obtained using martingale difference ideas and an elementary derivation is presented in Appendix A.
 Proof of Proposition 3.1..
@@
 $l_{n}$ $\doteq-\log_{2}\Pr\left(Y_{n}|{\mathbf{Y}}_{n-1}\right)\equiv-\log_{2}\left(p_{n}(Y_{n})\right)$ (15)
-We note that the random variables $Z_{n}\doteq l_{n}$ satisfy all of the conditions for Lemma B.2, if we impose the restriction that they are strictly bounded i.e.
+We note that the random variables $Z_{n}\doteq l_{n}$ satisfy all of the conditions for Lemma A.2, if we impose the restriction that they are strictly bounded i.e.
 $|l_{n}|<L$, for some some finite $L$.
@@
 In practice, this is easily satisfied by any practical LLM.
-Then from Lemma B.2, we have for any $\epsilon>0$:
+Then from Lemma A.2, we have for any $\epsilon>0$:
 $\lim_{N\rightarrow\infty}\Pr\left(\left|\frac{1}{N}\sum_{n=1}^{N}\left(l_{n}-H(p_{n})\right)\right|>\epsilon\right)$ $=0$ (16)
@@
 3.2 Informal Proof: Roads not Traveled
-We now present an alternative proof of Proposition 3.1 using a more informal argument.
-Specifically, we introduce the conceptual device of an auxiliary language model that will allow us to apply the traditional Law of Large Numbers to a sequence of dependent random variables.
-Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens.
+Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{Y}$ of three tokens.
 Also shown in red is the probability tree of an auxiliary model $M^{\prime}_{\mathbf{S}}$ for a string $\mathbf{S}$} with $S_{1}=3,~S_{2}=1$ and so on.
-3.2.1 Concept of an Auxiliary Language Model
-Let us consider a fixed string ${\mathbf{S}}$.
+Consider a fixed string $\mathbf{S}$}.
 For this particular string $\mathbf{S}$} and model M, we have a fixed set of distributions $q_{n}(y)\doteq\Pr(S_{n}=y|{\mathbf{S}}_{n-1})$.
@@
 The auxiliary model $M^{\prime}_{\mathbf{S}}$ is much simpler than the general language model M, because its output tokens $Y^{\prime}_{n}$ are generated independently of other tokens.
-Our informal proof of Proposition 3.1 is based on the following idea.
+Our informal proof of Proposition 3.1 is based on the following idea: while calculating $\Pr(E)$ for some event $E$, we can change any part of the probability tree that do not affect the event $E$ (“the roads not traveled”) to make our calculation easier e.g.
+by allowing us to work with an auxiliary language model $M^{\prime}_{\mathbf{S}}$ which generates strings of independent tokens.
+Informal Proof of Proposition 3.1..
+Assuming the tokens $Y_{n}$ are generated by the auxiliary model $M^{\prime}_{\mathbf{S}}$ (instead of the original model M), the random variables $l_{m},~l_{n}$ are independent and $\operatorname*{{E}}\left[l_{n}\right]\equiv H\left(q_{n}\right)$.
+Equation (16) then immediately follows from Proposition 2.2.
+∎
+This informal method allows us to treat the martingale property that the formal proof relies on as a quasi-independence condition to reason about the random variables $l_{n}$ intuitively.
+As an example, for the auxiliary model $M^{\prime}_{\mathbf{S}}$, from (12, 13), the variances of $l_{n},~l_{M}({\mathbf{S}}_{N})$ are respectively $\lambda^{2}(q_{n}),~\lambda^{2}_{M}({\mathbf{S}}_{N})$.
+We then have the following simple bound on deviations using the Chebyshev’s Inequality Cohen (2015):
+$\Pr\left(\left|l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right|>\alpha\lambda_{M}({\mathbf{S}}_{N})\right)$ $\leq\frac{1}{\alpha^{2}},~\forall\alpha>0$ (17)
+3.3 Typical Sets
+We can now formally define the typical set $\mathcal{T}_{M}^{n}(\epsilon)\subset\mathcal{Y}^{n}$ of typical strings ${\mathbf{Y}}_{n}$ for a model M with $\epsilon>0$:
+$\mathcal{T}_{M}^{n}(\epsilon)$ $\doteq\left\{{\mathbf{Y}}_{n}:\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}\equiv\left\{{\mathbf{Y}}_{n}:\left|h_{M}({\mathbf{Y}}_{n})+\frac{1}{n}\log_{2}\left(\Pr({\mathbf{Y}}_{n})\right)\right|<\epsilon\right\}$ (18)
+Proposition 3.1 asserts that with high probability a stochastic model $M$ is constrained to only output typical strings from this set $\mathcal{T}_{M}^{n}(\epsilon)$ for any $\epsilon>0$ if we consider sufficiently long strings.
+In practical applications of the typical set concept, we almost invariably find that the typical set is vanishingly small.
+This is what makes the typical set concept so powerful: it identifies the small number of outcomes that actually matter, so we can focus our resources on them efficiently.
+We will show that here too, the typical set must be vanishingly small under mild assumptions.
+Since the vast majority of token strings represent gibberish text, it is trivially true that any LLM that generates coherent text will not generate most token strings.
+Thus, the simple notion of smallness - that the typical set is a vanishing subset of all token strings - is not very interesting.
+We want to show a much stronger result that even if we exclude gibberish sentences and only look at the set of grammatically correct texts, any practical LLM can only generate a vanishingly small fraction of all possible grammatical texts.
+This requires additional analytical machinery which we now introduce.
+4 The Grammar Police
+Let $\mathcal{G}(n)$ denote a dictionary of “grammatical” strings ${\mathbf{Y}}_{n}$ of length $n$ and $g(n)\doteq\frac{1}{n}\log_{2}|\mathcal{G}(n)|$.
+Since most token strings represent meaningless gibberish, we expect $|\mathcal{G}(n)|\ll|\mathcal{Y}|^{n}$ or $g(n)\ll\log_{2}|\mathcal{Y}|$ for any reasonable notion of grammar.
+However, we also want our dictionary to be rich enough to include sentences of unlimited length and complexity to avoid trivialities.
+Formally we require that $g(n)\geq g_{min}>0$.
+Beyond this, our analysis below is very general, which means we can define the notion of “grammatical strings” to be essentially anything we want.
+4.1 Augmented Typical Sets for Language Models
+We can now define the augmented typical set of the model $M$ for dictionary $\mathcal{G}(n)$ as:
+$\mathcal{T}_{M,G}^{n}(\epsilon)$ $\doteq\mathcal{G}(n)\cap\mathcal{T}_{M}^{n}(\epsilon)\equiv\left\{{\mathbf{Y}}_{n}\in\mathcal{G}(n):\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}$ (19)
+Let $p_{G}(n)\doteq\Pr({\mathbf{Y}}_{n}\in\mathcal{G}(n))$ be the probability that model $M$ generates a grammatically correct string ${\mathbf{Y}}_{n}\in\mathcal{G}(n)$.
+We will assume that $p_{G}(n)\geq p_{G}>0$ i.e.
+the model generates grammatical strings of any length with a non-zero probability.
+Informally, the model $M$ knows how to generate grammatically correct sentences when suitably prompted, but we do not require it to be perfect and it may still sometimes generate ungrammatical outputs.
+The probability that the model $M$ generates a grammatical string ${\mathbf{Y}}_{n}\in\mathcal{G}(n)$ can be written as:
+$\Pr\big{(}{\mathbf{Y}}_{n}\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}$ $=\frac{\Pr\big{(}{\mathbf{Y}}_{n}\big{)}}{p_{G}(n)}\equiv\frac{2^{-nl_{M}({\mathbf{Y}}_{n})}}{p_{G}(n)}$ (20)
+We can use (20) to relate the perplexity of grammatical strings specifically to the entropy analogous to (11):
+$\frac{1}{n}H\left(\Pr({\mathbf{Y}}_{n})\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\right)$ $\equiv\frac{\log_{2}p_{G}(n)}{n}+\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{n})\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\right]\leq g(n)$ (21)
+where the last inequality holds with equality for a model that selects all grammatical texts with equal probability: $\Pr({\mathbf{Y}}_{n})\equiv\frac{1}{|\mathcal{G}(n)|},~\forall{\mathbf{Y}}_{n}\in\mathcal{G}(n)$ (for which $p_{G}(n)=1$).
+Unlike (11), this does not tell us anything about the empirical entropy $h_{M}({\mathbf{Y}}_{n})$ of individual strings ${\mathbf{Y}}_{n}$.
+We will now show that we can safely ignore high entropy strings.
 Proposition 4.1.
-The Indifference Principle.
-For any random event $E$ defined entirely on the subtree of a particular string ${\mathbf{S}}$, $\Pr(E)$ will be the same for generative model M as for the auxiliary model $M^{\prime}_{\mathbf{S}}$.
-The idea behind Proposition 3.2 is that while calculating $\Pr(E)$ for some event $E$, we can change any part of the probability tree that do not affect the event $E$ (“the roads not traveled”) to make our calculation easier e.g.
-by allowing us to work with an auxiliary language model $M^{\prime}_{\mathbf{S}}$ which generates strings of independent tokens.
-For our purposes, Proposition 3.2 applies to all events in the $\sigma$-algebra generated by the conditional likelihood random variables $l_{n}$ as defined in (16).
-Informal Proof of Proposition 3.1..
-Using the Proposition 3.2, we can evaluate the distribution of random variables $l_{n}$ assuming the tokens $Y_{n}$ are generated by the auxiliary model $M^{\prime}_{\mathbf{S}}$ instead of the original model M.
-Then the random variables $l_{m},~{}l_{n}$ are independent of each other and $\operatorname*{{E}}\left[l_{n}\right]\equiv H\left(q_{n}\right)$.
-The weak Law of Large Numbers for the sequence of independent zero-mean random variables $l_{n}-H\left(q_{n}\right)$ immediately gives us (17).
-∎
-Proposition 3.2 allows us to use traditional statistical laws that apply to sequences of independent random variables directly without having to prove bespoke versions of those laws such as Proposition B.1.
-For instance, we may be able to use Proposition 3.2 to formulate a version of the Central Limit Theorem for $\frac{1}{\lambda_{M}({\mathbf{S}}_{N})}\left(l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right)$.
-We delegate a full development of these ideas to future work and limit ourselves to one simple non-asymptotic result to illustrate the method.
-3.3 Chebyshev Bound on Deviations
-We note that for the auxiliary model $M^{\prime}_{\mathbf{S}}$, the variance of the conditional log likelihood random variable $l_{n}$ is $\lambda^{2}(q_{n})$ as defined in (13).
-Assuming independence of $l_{n}$ using Proposition 3.2, we can show that the variance of $l_{M}({\mathbf{S}}_{N})$ is $\lambda^{2}_{M}({\mathbf{S}}_{N})$ as defined in (14).
-We can then write using the Chebyshev’s Inequality (Cohen, 2015) for any $\alpha>0$:
-$\Pr\left(\left|l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right|>\alpha\lambda_{M}({\mathbf{S}}_{N})\right)$ $\leq\frac{1}{\alpha^{2}}$ (18)
-3.4 Typical Sets are Vanishingly Small
-Let us now formally define the typical set $\mathcal{P}_{M}^{n}(\epsilon)\subset\mathcal{T}^{n}$ of typical strings ${\mathbf{Y}}_{n}$ for a model M with $\epsilon>0$:
-$\mathcal{P}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}$ (19)
-$\equiv\left\{{\mathbf{Y}}_{n}:\left|h_{M}({\mathbf{Y}}_{n})+\frac{1}{n}\log_{2}\left(\Pr({\mathbf{Y}}_{n})\right)\right|<\epsilon\right\}$
-Proposition 3.1 asserts that with high probability a stochastic model M is constrained to only output typical strings from this set $\mathcal{P}_{M}^{n}(\epsilon)$ for any $\epsilon>0$ if we consider sufficiently long strings.
-This observation has profound consequences.
-We will now show one such consequence, namely that the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ is very small:
+Entropies cannot be too large.
+The empirical entropy $h_{M}({\mathbf{Y}}_{n})$ of any long grammatical string ${\mathbf{Y}}_{n}\in\mathcal{G}(n)$ generated by the model $M$ is greater than $g(n)\equiv\frac{1}{n}\log_{2}|\mathcal{G}(n)|$ with vanishing probability:
+$\lim_{n\rightarrow\infty}$ $\Pr\big{(}h_{M}({\mathbf{Y}}_{n})>g(n)+\epsilon\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}=0,~\forall\epsilon>0$ (22)
+Proof.
+First we note that for sufficiently large $n$, we must have with high probability $\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\frac{\epsilon}{2}$. Then we have:
+$\Pr\big{(}h_{M}({\mathbf{Y}}_{n})>g(n)+\epsilon\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}$ $\leq\frac{1}{p_{G}(n)}\Pr\big{(}\mathcal{E}_{1}(n)\big{)}\leq\frac{1}{p_{G}}\Pr\big{(}\mathcal{E}_{1}(n)\big{)}$
+where $\mathcal{E}_{1}(n)\doteq\left\{{\mathbf{Y}}_{n}\in\mathcal{G}(n):l_{M}({\mathbf{Y}}_{n})>g(n)+\frac{\epsilon}{2}\right\}$. We can then bound this probability as:
+$\Pr(\mathcal{E}_{1}(n))$ $=\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{1}(n)}\Pr({\mathbf{Y}}_{n})\leq\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{1}(n)}2^{-n(g(n)+\frac{\epsilon}{2})}\equiv\left|\mathcal{E}_{1}(n)\right|2^{-n(g(n)+\frac{\epsilon}{2})}<\left|\mathcal{G}(n)\right|2^{-n(g(n)+\frac{\epsilon}{2})}\equiv 2^{-\frac{n\epsilon}{2}}$
+Thus we have shown that $\Pr\big{(}h_{M}({\mathbf{Y}}_{n})>g(n)+\epsilon\big{|}{\mathbf{Y}}_{n}\in\mathcal{G}(n)\big{)}<\frac{1}{p_{G}}2^{-\frac{n\epsilon}{2}}\rightarrow 0$ as $n\rightarrow\infty$. ∎
+Proposition 4.1 allows us to purge high entropy strings out of the typical set while retaining all of its probability mass.
+Specifically, we define the purged typical set $\mathcal{P}_{M,G}^{n}(\epsilon)$ as:
+$\mathcal{P}_{M,G}^{n}(\epsilon)$ $\doteq\mathcal{G}(n)\cap\mathcal{T}_{M}^{n}(\epsilon)\cap\mathcal{E}_{1}^{c}(n),~\mathrm{where}~\lim_{n\rightarrow\infty}\Pr\left(\mathcal{P}_{M,G}^{n}(\epsilon)|{\mathbf{Y}}_{n}\in\mathcal{G}(n)\right)=1,~\forall\epsilon>0$ (23)
+4.2 The Essential Smallness of Typical Sets
+For an entropy-maximizing model that chooses all grammatical strings with equal probability, the typical set includes all grammatical strings $\mathcal{P}_{M,G}^{n}(\epsilon)\equiv\mathcal{G}(n)$.
+We now formalize the assumption that practical LLMs do not do this, and for all such models, we will show that the typical set is a vanishingly small subset of $\mathcal{G}(n)$.
+Assumption 4.1.
+Practical LLMs are not entropy-maximizing.
+The empirical entropy $h_{M}({\mathbf{Y}}_{n})$ is strictly smaller than $g(n)$ for at least some fraction of grammatical typical strings.
+Specifically, there is an entropy gap $\Delta g>0$ such that $h_{M}({\mathbf{Y}}_{n})\leq g(n)-\Delta g$ for at least some non-zero fraction $\rho>0$ of typical strings ${\mathbf{Y}}_{n}$:
+$\left|\mathcal{E}_{2}(n)\right|\geq\rho\left|\mathcal{P}_{M,G}^{n}(\epsilon)\right|,~\mathrm{where}~\mathcal{E}_{2}(n)\doteq\left\{{\mathbf{Y}}_{n}\in\mathcal{P}_{M,G}^{n}(\epsilon):h_{M}({\mathbf{Y}}_{n})\leq g(n)-\Delta g\right\}$ (24)
 Proposition 4.2.
-The size of the typical set $\mathcal{P}_{M}^{n}$ is an exponentially vanishing fraction $02^{-\gamma n},~{}\gamma>$ 0 of all grammatically correct sequences.
-The argument below closely tracks standard information-theoretic results (Cover & Thomas, 2005a) about the size of the typical set. We start by restating Proposition B.1 informally as:
-$\Pr({\mathbf{Y}}_{n})$ $\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\rightarrow 2^{-nh_{M}({\mathbf{Y}}_{n})}$ (20)
-Let $T_{g}(n)$ be the total number of grammatically correct strings ${\mathbf{Y}}_{n}$ of length $n$.
-The average entropy $h_{M}({\mathbf{Y}})$ is a measure of the richness of the model that generated the string.
-Language models typically have a “temperature” parameter $T$ in the range $T\in[0,1]$ that can be used to tune the richness of the model’s output.
-Higher temperature means less “predictable” (i.e.
-less deterministic) and more “creative” (i.e.
-more random) outputs.
-This corresponds to higher values for $h_{M}({\mathbf{Y}})$.
-Consider the maximally rich model of length $n$ that has the highest possible entropy $h_{M}({\mathbf{Y}}_{n})\equiv h_{max}(n)$ which corresponds to choosing all possible (grammatically correct) strings of length $n$ with equal probability [^3] Note that this does not mean that the token distributions $p_{m}(y)$ are uniform over all valid tokens at each step $m$; indeed tokens corresponding to richer sub-trees as shown in Fig.
-1 should be chosen with higher probability to maximize the overall richness of the model’s output..
-By definition, for the maximally rich model, the probability of choosing any valid string is $2^{-nh_{max}(n)}$ and this means the number of possible strings of length $n$ is:
-$T_{g}(n)$ $\equiv 2^{nh_{max}(n)}$ (21)
-We can take (21) as the defining relation for $h_{max}(n)$.
-In practical LLMs, even at the highest temperature setting i.e.
-$T=1$, the token probability is typically concentrated in a small number of tokens and the token probabilities are determined by far more complex rules than the richness of their sub-trees.
-Let us formalize this as an assumption.
-Assumption 3.1.
-Practical LLMs are not entropy-maximizing. Any language model M has an entropy gap $\Delta h>0$ such that the following holds for all sufficiently long strings:
-$h_{M}({\mathbf{Y}}_{n})$ $<h_{max}(n)-\Delta h,~{}\forall{\mathbf{Y}}_{n}\in\mathcal{T}^{n}$ (22)
-Proof of Proposition 3.3.
-From (20), for $n\gg 1$, we have $\Pr({\mathbf{Y}}_{n})\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\to 2^{-nh_{M}({\mathbf{Y}}_{n})}>2^{-n(h_{max}(n)-\Delta h)}$ where we used Assumption 3.1 in the last step.
-Since this holds for any typical string of length $n$, we must have for small $\epsilon>0$ and correspondingly large $n$:
-$\left|\mathcal{P}_{M}^{n}(\epsilon)\right|$ $<2^{n\left(h_{max}(n)-\Delta h)\right)}\equiv 2^{-n\Delta h}\left|T_{g}(n)\right|$ (23)
-Thus, the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ represents an exponentially vanishing fraction of all grammatically correct strings. ∎
-We emphasize that all we needed to establish Proposition 3.3 was Assumption 3.1 which only requires that the model assigns higher probability to some sentences than to others.
-3.5 Over-Typical Sets are Vanishingly Smaller Still
+Typical sets are small. Under Assumption 4.1, the size of the pruned typical set $\mathcal{P}_{M,G}^{n}$ is an exponentially vanishing fraction of all grammatically correct sequences $\mathcal{G}(n)$.
+Proof.
+From Proposition 3.1, for sufficiently large $n$, we have $|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})|\leq\frac{\Delta g}{2}$.
+Using this observation for ${\mathbf{Y}}_{n}\in\mathcal{E}_{2}(n)$, we have $l_{M}({\mathbf{Y}}_{n})\leq g(n)-\frac{\Delta g}{2}$.
+Then we have:
+$1$ $\geq\Pr(\mathcal{E}_{2}(n))\equiv\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{2}(n)}2^{-nl_{M}({\mathbf{Y}}_{n})}\geq\sum_{{\mathbf{Y}}_{n}\in\mathcal{E}_{2}(n)}2^{-n\left(g(n)-\frac{\Delta g}{2}\right)}\equiv\frac{|\mathcal{E}_{2}(n)|}{|\mathcal{G}(n)|}2^{\frac{n\Delta g}{2}}\geq\frac{\rho|\mathcal{P}_{M,G}^{n}(\epsilon)|}{|\mathcal{G}(n)|}2^{\frac{n\Delta g}{2}}$ (25)
+Thus, we have shown $|\mathcal{P}_{M,G}^{n}(\epsilon)|\leq|\mathcal{G}(n)|2^{-n\left(\frac{\Delta g}{2}-\frac{\log_{2}(\rho)}{n}\right)}$. ∎
 The obverse of Proposition 4.2 is that almost all grammatically correct strings are non-typical for a given language model M.
@@
 Formally, we define the set $\mathcal{V}_{M}^{n}(\epsilon)$ of over-typical strings of length $n$ for a model M as:
-$\mathcal{V}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left(h_{M}({\mathbf{Y}}_{n})-l_{M}({\mathbf{Y}}_{n})\right)>\epsilon\right\}$
-$\equiv\left\{{\mathbf{Y}}_{n}:\Pr({\mathbf{Y}}_{n})>2^{-n\left(h_{M}({\mathbf{Y}}_{n})-\epsilon\right)}\right\}$ (24)
+$\mathcal{V}_{M}^{n}(\epsilon)$ $\doteq\left\{{\mathbf{Y}}_{n}\in\mathcal{G}(n):l_{M}({\mathbf{Y}}_{n})\leq h_{M}({\mathbf{Y}}_{n})-\epsilon<g(n)-\epsilon\right\}$ (26)
 Proposition 4.3.
-The size of the over-typical set $\mathcal{V}_{M}^{n}(\epsilon)$ is an exponentially vanishing fraction of all grammatically correct sequences.
+Over-Typical sets are smaller still. The size of the over-typical set $\mathcal{V}_{M}^{n}(\epsilon)$ is an exponentially vanishing fraction of all grammatically correct sequences.
 Proof.
-For any ${\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)$, we have from (24) and (22):
-$\Pr({\mathbf{Y}}_{n})$ $>2^{-nh_{max}(n)}\times 2^{n(\Delta h+\epsilon)}$ (25)
-Therefore, we must have
-$\left|\mathcal{V}_{M}^{n}(\epsilon)\right|$ $<\left|T_{g}(n)\right|2^{-n(\Delta h+\epsilon)}$
-where we used the definition of $h_{max}$ in (21). ∎
-We have shown that almost all long grammatically correct strings are under-typical for any language model M. We illustrate these ideas for a simple Markov language model in Appendix A.
+The argument is similar to the proof of Proposition 4.2:
+$1$ $\geq\Pr\left({\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)\right)\equiv\sum_{{\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)}2^{-nl_{M}({\mathbf{Y}}_{n})}\geq\left|\mathcal{V}_{M}^{n}(\epsilon)\right|2^{-n(g(n)-\epsilon)}\equiv\frac{\left|\mathcal{V}_{M}^{n}(\epsilon)\right|}{\left|\mathcal{G}(n)\right|}2^{n\epsilon}$ (27)
+Thus we have $\left|\mathcal{V}_{M}^{n}(\epsilon)\right|\leq\left|\mathcal{G}(n)\right|2^{-n\epsilon}$ which proves the result. ∎
+For Assumption 4.1 to be violated, we need the perplexities of almost all typical strings to be almost equal; conversely, any model that chooses some strings with higher probabilities than others must satisfy Assumption 4.1 and therefore Proposition 4.2.
+Together with Proposition 4.3, this means that almost all grammatical strings are undertypical for any practical LLM.
+4.3 Constriction by Staying True to Type
+We now present some example consequences of Proposition 4.2 as a payoff for the preceding analysis.
+Informally, Proposition 3.1 says that all models must stay “true to type”, and Proposition 4.2 says that this narrowly circumscribes the range of a model’s possible behaviors and outputs.
+Consider the augmented language model $G=(M,D)$ shown in Fig.
+2, where the output of a model $M$ is screened by a grammar-checking device $D$ which defines the dictionary $\mathcal{G}(n)$; if the model’s output $\mathbf{Y}$} is not grammatically correct, then it is prompted again to produce a new random string $\mathbf{Y}$} until it generates a grammatical string [^3] Note that we assume that the device $D$ is deterministic, which means our analysis does not allow the grammar checking device to be another stochastic LLM..
+The construction in Fig.
+2 implicitly assumes a decidable grammar, which means it describes a “recursive language” in Type 1 of the Chomsky hierarchy - the same category natural languages are believed to belong to (Chomsky, 2014).
+Figure 2: Language Model Augmented with a Grammar Checking Device.
+Since $p_{G}(n)>p_{G}>0$, the augmented model $G$ is guaranteed to generate a grammatical string in a finite time with probability one and the probability that model $G$ generates a grammatical string ${\mathbf{Y}}_{n}$ is $\Pr({\mathbf{Y}}_{n}|{\mathbf{Y}}_{n}\in\mathcal{G}(n))$.
+Furthermore, we see from (23) that the augmented model $G$ will only generate outputs from the set $\mathcal{P}_{M,G}^{n}(\epsilon)$ with high probability.
+We can then interpret the Proposition 4.2 as a restriction on what the augmented model $G$ can generate as illustrated in the following examples:
+  1. Ex 1
+    The Python programming language.
+We define $\mathcal{G}(n)$ as the set of all syntactically correct Python code that compiles (or interprets) without error.
+While we know LLMs are capable of generating complex Python code, Proposition 4.2 asserts that any model is limited to generating a vanishing fraction of all valid Python programs.
+  2. Ex 2
+    Proofs in First Order Logic (FOL).
+We define $\mathcal{G}(n)$ as the set of valid mathematical proofs in FOL.
+A model used to generate proofs in FOL is limited to generating a vanishing fraction of all correct proofs.
+  3. Ex 3
+    Lyrical Poetry. We define $\mathcal{G}(n)$ as the set of stories consisting of a sequence of haikus. A model is limited to a vanishing fraction of all such possible stories.
 5 Experiments with Open-Source Models
-We performed a series of experiments to verify the ideas described in Section 3.
+We performed a series of experiments to verify the ideas described in Sections 3 and 4.
 The basic idea is to use a local instance of a pre-trained open-source model to evaluate the empirical entropy of a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions.
-We ran experiments on the smaller GPT-2 model (Radford et al., 2019) as well as the more recent, larger and more capable Llama 3.1 8B (Touvron et al., 2023) model.
+We used the smaller GPT-2 model (Radford et al., 2019) as well as the larger and more capable Llama 3.1 8B (Touvron et al., 2023) model.
 5.1 Perplexity of Self-Generated Strings
@@
 We want the number of tokens $N_{max}$ to be large enough to observe the asymptotic behavior of the perplexity of the generated string.
-In practice, $N_{max}$ is constrained by certain limitations of the models as we discuss in the sequel.
 The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing.
-We ran this experiment many times for different seed values for the random number generator, different choices of sampling distributions and a number of different initial prompts.
-For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3.1.
+We ran this experiment many times for different seed values for the random number generator, different choices of sampling distributions and initial prompts.
+For each generated string, we can then test our claim in Proposition 3.1.
 Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
@@
 We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 5.1.
-However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that our model will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^4] Note that our choice of top-k sampling is a compromise.
+However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that our model will output the actual next token $Y_{n}$ from the string $\mathbf{Y}$} which may not be included in the list of top-$k$ tokens [^4] Our choice of top-k sampling is a compromise.
 The number of tokens in GPT-2 or Llama is quite large ($|\mathcal{T}|\approx 40,000$).
-If we exhaustively list the full probability distribution of all tokens, our experiment becomes very computationally expensive and slow..
+If we exhaustively list the full probability distribution of all tokens, the experiment becomes very computationally expensive and slow..
 When this happens, we replaced the lowest probability entry in the list of top-k tokens with the actual next token from our string along with its probability.
 The normalized probability of each of the top-k tokens along with their token-ID and token strings, as well as the probability and ID of the actual next token are saved in a json file.
-This json file is then processed in exactly the same way as in Section 4.1 to compute the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the sub-string ${\mathbf{Y}}_{N}$ for the GPT-2 model for $N=1\dots N_{max}$.
+This json file is then processed in exactly the same way as in Section 5.1 to compute $l_{M}({\mathbf{Y}}_{N}),~h_{M}({\mathbf{Y}}_{N})$ and $\lambda_{M}({\mathbf{Y}}_{N})$ of the sub-strings ${\mathbf{Y}}_{N},~N=1\dots N_{max}$ for the GPT-2 or Llama 3.1 8B model.
 5.3 Results and Discussion
+(a) GPT-2, $N_{max}=512$.
+(b) Llama 3.1 8B, $N_{max}=10240$.
+Figure 3: Log-perplexity & empirical entropy for GPT-2 or Llama 3.1, self-generated text, top-100 sampling.
 Figure 3 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 or Llama 3.1 8B model as a function of $N$ for several text strings generated by the same model.
 The plots also shows the range $h_{M}({\mathbf{Y}}_{N})\pm\lambda_{M}({\mathbf{Y}}_{N})$ and $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$.
-We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently in all cases.
-The asymptotic convergence is more clearly seen in the long strings from the Llama 3.1 8B model [^5] The GPT-2 model cannot generate similarly long texts.
+We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently in both cases.
+The asymptotic convergence is more clearly seen in the long string from the Llama 3.1 8B model [^5] The GPT-2 model cannot generate similarly long texts.
 in Fig.
-2(c), 2(d) .
-We find that very long synthetic texts such as those in Fig.
-2(c), 2(d) can be ungrammatical, repetitive and generally of poor quality, as reported elsewhere (Holtzman et al., 2020).
-However, Proposition 3.1 holds regardless of the quality of the text as seen in Fig.
-2.
-(a) GPT-2, $N_{max}=512$.
-(b) Llama 3.1 8B, $N_{max}=1024$.
-(c) Llama 3.1 8B, $N_{max}=10240$.
-(d) Llama 3.1 8B, $N_{max}=10240$.
-Figure 2: Log-perplexity and empirical entropy for GPT-2 or Llama 3.1 8B of self-generated text, top-100 sampling.
-Figure 3 shows the log-perplexity and empirical entropy for the Llama 3.1 8B model for two long, non-Llama generated strings.
+3(b).
+Additional plots showing the asymptotic behavior are presented in Appendix LABEL:sec:more_exp.
+However, we find that very long synthetic texts such as those in Fig.
+3(b) can be ungrammatical and generally of poor quality, as reported elsewhere Holtzman et al.
+(2020).
+While Proposition 3.1 holds regardless of the quality of the text, its practical relevance depends on whether the predicted asymptotic behavior is observable in normal-length texts.
+Figure 4 shows the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ against the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ for the Llama-3.1 8B model for several text strings generated by the same model.
+A total of $450$ strings were generated using top-$100$ sampling from $15$ distinct prompts and $30$ different random seeds for each prompt.
+Of these, abnormally short strings of less than $100$ tokens were excluded from the analysis.
+The remaining $383$ strings had an average length of $685$ tokens and a standard deviation of $321$.
+The clustering of the log-perplexity around the $45^{\circ}$ line in Fig.
+4 provides direct visual affirmation for Proposition 3.1.
+Figure 4: Log-perplexity & empirical entropy of “normal-length” text generated by Llama-3.1 8B.
+The fact that the perplexity of the generated strings vary over a wide range in Fig.
+4 also provides direct evidence for the Assumption 4.1 that the Llama-3.1 8B model is not entropy-maximizing and therefore must have vanishingly small typical sets per Proposition 4.2.
+Additional evidence for Propositions 4.2 and 4.3 is in Fig.
+5 that analyses two long, non-Llama generated strings.
 The text used in Fig.
-3(a) is an excerpt from a recent article (Pelly, 2025) in Harper’s Magazine, a respected literary publication.
-The string used in Fig.
-3(b) was generated by ChatGPT o1 Pro (OpenAI, 2024).
+5(a) is an excerpt from a recent article Pelly (2025) in Harper’s Magazine, a respected literary publication.
+The text in Fig.
+5(b) was generated by ChatGPT o1 Pro OpenAI (2024).
+We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ for Llama 3.1 8B is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ larger than the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ for both strings in Fig.
+5 showing that these strings are clearly under-typical for the Llama 3.1 8B model.
+If we assume that the strings in Fig.
+5 are randomly sampled from the set of all grammatically correct texts, then this under-typicality is a simple consequence of the smallness of the Llama-3.1 8B model’s typical sets.
 (a) Human writing.
@@
 Figure 5: Log-perplexity and empirical entropy for Llama 3.1 8B of non-Llama-generated strings $\mathbf{Y}$.
-We see from Fig.
-3(a) and Fig.
-3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ for Llama 3.1 8B is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ larger than the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ showing that the corresponding strings are clearly under-typical for the Llama 3.1 8B model.
-While this non-convergence is not required by our theory - we have not considered whether the converse to Proposition 3.1 is true - it is strongly suggested by Propositions 3.3 and 3.4 if we assume that the strings in Fig.
-3 are randomly sampled from the set of all grammatically correct texts.
 5.4 More Catholic than the Pope
+This brings us to the remarkable plots in Fig.
+6 which shows the log-perplexity of excerpts from classic texts for Llama 3.1 8B, specifically from Alice in Wonderland and The Great Gatsby in Figs.
+6(a) and 6(b) respectively.
+The average entropy for the token distributions in Fig.
+6(a) is an astonishingly small $0.11$ bits, which means that the model on average assigns a very high probability ($\sim 99\%$) to the precise sequence of tokens from the excerpt for its predicted next tokens.
+In other words, this excerpt has been memorized by the Llama 3.1 8B model in the sense of eidetic memorization Carlini et al.
+(2021): with zero temperature, it will output this text verbatim.
+The same observation also applies to Fig.
+6(b).
+We also see from both plots that the log-perplexity is several standard deviations below the empirical entropy making the excerpts slightly, but unambiguously, over-typical for the Llama-8B model: the model judges these texts to be more perfect than its own creations!
+From Proposition 4.3, we can conclude that these excerpts are among the small number of very special strings to be so memorized by the model.
+Clearly, these texts were part of the training set of the Llama-8B model and they were esteemed highly by its cost function.
 (a) Alice in Wonderland excerpt.
@@
 Figure 6: Log-perplexity and empirical entropy for Llama 3.1 8B of excerpts from literary classics.
-This brings us to the remarkable plots in Fig.
-4 which shows the log-perplexity of excerpts from classic texts for Llama 3.1 8B, specifically from Alice in Wonderland and The Great Gatsby in Figs.
-4(a) and 4(b) respectively.
-The average entropy for the token distributions in Fig.
-4(a) is an astonishingly small $0.11$ bits, which means that the model on average assigns a very high probability ($\sim 99\%$) to the precise sequence of tokens from the excerpt for its predicted next tokens.
-In other words, this excerpt has been memorized by the Llama 3.1 8B model.
-The same observation also applies to Fig.
-4(b).
-We also see from both plots that the log-perplexity is several standard deviations below the empirical entropy making the excerpts slightly, but unambiguously, over-typical for the Llama-8B model: the model judges these texts to be more perfect than its own creations!
-From Proposition 3.4, we can conclude that these excerpts are among the small number of very special strings to be so memorized by the model.
-Clearly, these texts were part of the training set of the Llama-8B model and they were esteemed highly by its cost function.
-Interestingly, the over-typicality of these excerpts means that it is pretty much impossible for the Llama-8B model itself to generate them verbatim - even with a $99\%$ probability that each next token will exactly match the excerpt, a randomly generated string can be expected to diverge from the excerpt after only about $100$ tokens.
-Additional plots and more details about these experiments are presented in Appendix C.1.
-5 Practical Applications
-Consider the problem of AI text detection i.e.
-determining whether a string ${\mathbf{Y}}$ was generated by a model M.
-The intuition behind (9) suggests a simple method for doing this: check if the perplexity is small enough.
-Specifically, declare ‘yes’ if $l_{M}({\mathbf{Y}})<l_{th}$ for a threshold $l_{th}$.
-The same test can also be used for the problem of LLM dataset inference i.e.
-determining whether a string ${\mathbf{Y}}$ was used as part of the training dataset for the model M.
-While, this idea is not new and indeed variations of it are behind popular solutions for both problems (Mitchell et al., 2023; Maini et al., 2024), our results in this paper provide a rigorous theoretical foundation for it.
-It also provides a principled way to choose the classification threshold $l_{th}$ which previously had to be computed empirically (Carlini et al., 2021).
-As an example, with $l_{th}=h_{M}({\mathbf{Y}})+3\lambda_{M}({\mathbf{Y}})$, Fig.
+In summary, we see from Figs.
+3 and 4 that synthetic texts generated by an LLM are typical for the model, while most external texts such as in Fig.
+5 are undertypical.
+By definition, over-typical strings are very special such as the literary classics in Fig.
+6 that have been memorized by the model.
+6 Practical applications: the capybara problem
+A simple and direct application of the ideas in this work is for testing whether a given piece of text has “abnormally” high or low perplexity for a language model.
+Such a test is relevant to many important practical problems.
+Low perplexity is commonly used as one marker of synthetic text in AI text detection tools (Wu et al., 2025) by testing $l_{M}({\mathbf{Y}}_{n})\begin{subarray}{c}<\\ >\end{subarray}l_{th}$ against a threshold $l_{th}$.
+However, it turns out that choosing a suitable classification threshold $l_{th}$ is difficult - the so-called “capybara problem” (Hans et al., 2024).
+Our theory shows a principled way to do this: with $l_{th}=h_{M}({\mathbf{Y}}_{n})+3\lambda_{M}({\mathbf{Y}}_{n})$, Fig.
 4 shows that we can accurately distinguish between synthetic text generated by Llama 3.1 8B and non-Llama text such as in Fig.
-3 even with relatively short strings of $1000$ tokens.
-We can also provide performance guarantees.
-Indeed, a simple lower-bound on the probability of a false negative (i.e.
+5 even with relatively short strings of $\approx 500$ tokens.
+The key insight is to choose variable rather than fixed thresholds, i.e.
+different $l_{th}$ for different strings ${\mathbf{Y}}_{n}$.
+Our theory can also provide some performance guarantees: e.g.
+a simple upper-bound on the false negative rate (i.e.
 failing to detect a Llama-generated string) can be obtained from (17).
-Similar comments apply to the dataset inference problem: while it is well-known that text strings memorized by an LLM have low perplexity, our theory allows us to rigorously derive threshold values and performance guarantees.
+Note, however, that we cannot provide bounds on false positive because the converse to Proposition 3.1 does not hold in general without additional assumptions.
+Similar comments apply to the dataset inference problem (Maini et al., 2024): while it is well-known (Carlini et al., 2021) that text strings memorized by an LLM have low perplexity, our theory allows us to rigorously derive threshold values and performance guarantees.
 Extrapolating from Fig.
-4, we posit that if a long excerpt from a human-written book is over-typical for a language model, it is a near-statistical certainty that the book was part of the training data set for that model.
+6, we posit that if a long excerpt from a human-written book is over-typical for a language model, it is a near-statistical certainty that the book was part of the training data set for that model.
 7 Conclusions
 We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property.
-While this paper focused on a rigorous exposition of the AEP property, it opens up many questions for further study.
-An important open problem is to design practical algorithms for AI text detection and dataset inference using the typical set concept.
-Tighter non-asymptotic bounds on deviations that improve on (18) will significantly benefit practical applications by providing better performance guarantees.
-More fundamentally, Proposition 3.3 leads naturally to the question of how much do the typical sets of different models overlap.
-The theory presented in this paper is neutral on this question which must be answered empirically and both possible answers have interesting implications.
-No overlap between typical sets means that each model can be associated with a unique statistical signature.
-On the other hand, independently trained models converging on a common set of typical strings, would be a notable linguistic phenomenon, though it is hard to speculate on what this might mean.
-The AEP is also just one example of the broader observation that generative AI models are stochastic machines subject to statistical laws and the discovery of other statistical regularities in the outputs of generative models is an important topic for future work.
-Appendix A A Simple Markov Language Model
-We now illustrate the ideas from Section 3 for a simple Markov language model, $M_{1}$ with a binary token space consisting of only two tokens $\mathcal{T}=\{t_{1},t_{2}\}$ and a parameter $\rho\in(0,1)$.
-The user prompt ${\mathbf{X}}$ consists of a single token which we can set to be $t_{1}$ without loss of generality: ${\mathbf{X}}=[t_{1}]$.
-The model $M_{1}$ generates a text string of length $N$ following a simple probabilistic algorithm: the $n$’th token $Y_{n}$ is set equal to the $n-1$’th token with probability $\rho$.
-The prompt serves as the initial token i.e.
-$Y_{0}\doteq t_{1}$.
-Let $h_{b}(\rho)\doteq-\rho\log_{2}\rho-(1-\rho)\log_{2}(1-\rho)$ denote the binary entropy function with argument $\rho$.
-It is easy to see that $H(p_{n})\equiv h_{b}(\rho),~{}\forall n$ and $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(\rho)$ for any text string ${\mathbf{Y}}$.
-For $\rho=0.9$, $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(0.9)\approx 0.47$
-Example A.1.
-Consider the string ${\mathbf{Y}}_{1}$ with all tokens $Y_{n}=t_{1},~{}\forall n=1\dots N$.
-The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(\rho)$.
-For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(0.9)\approx 0.15$.
-Example A.2.
-Consider the string ${\mathbf{Y}}_{2}$ with alternating token values: ${\mathbf{Y}}_{2}\equiv[t_{2},t_{1},t_{2},t_{1},t_{2},t_{1}\dots]$.
-The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(1-\rho)$.
-For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(0.1)\approx 3.3$.
-Note that ${\mathbf{Y}}_{1}$ and ${\mathbf{Y}}_{2}$ are the most likely and least likely string of length $N$ respectively for $\rho=0.9$ (or indeed for any $\rho>0.5$).
-However, it is clear that neither string is typical for any reasonable $\epsilon$ with the perplexity of ${\mathbf{Y}}_{1}$ being abnormally low and that of ${\mathbf{Y}}_{2}$ abnormally high respectively.
-Example A.3.
-Consider the string ${\mathbf{Y}}_{3}$ created by flipping a fair coin repeatedly to generate a vector whose elements are iid randomly selected with $\Pr(t_{1})=\Pr(t_{2})=0.5$.
-Since $Y_{3}$ is a random string, its perplexity $l_{M_{1}}({\mathbf{Y}}_{3})$ is also a random variable.
-However, for long strings i.e.
-$N\gg 1$, with high probability, $l_{M_{1}}({\mathbf{Y}}_{3})\approx-0.5\log_{2}\rho-0.5\log_{2}(1-\rho)$.
-For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{3})\approx 1.74$ which is incidentally the arithmetic mean of $l_{M_{1}}({\mathbf{Y}}_{1})$ and $l_{M_{1}}({\mathbf{Y}}_{2})$.
-Clearly, with high probability, the string ${\mathbf{Y}}_{3}$ is not typical.
-Indeed, from these examples, we can see that typical strings for the model $M_{1}$ are precisely those that contain close to $(1-\rho)N$ token ’flips’ i.e.
-instances when $Y_{n+1}\neq Y_{n}$.
-Appendix B A Generalized Law of Large Numbers for Random Variables with Arbitrary Dependence
+We used this property to extend the information theoretic concept of typical set to language models.
+The main take-aways are as follows:
+  1. 1.
+    The classical AEP and the concept of typical sets can be generalized to apply to the outputs of LLMs.
+  2. 2.
+    These typical sets can be shown to be a vanishingly small subset of possible ”good” outputs under very general conditions and this means that LLMs are strongly constrained in what they can generate.
+  3. 3.
+    Experimentally we see that the asymptotic behavior in the AEP can be observed for LLMs for moderate length texts (a few hundred tokens), and thus may be practically relevant to detecting texts of abnormally high or low perplexity.
+  4. 4.
+    Synthetic texts from LLMs are random processes with known statistics (unlike natural text), and discovering other laws (e.g.
+Central Limit Theorems) that apply to LLM outputs is an open research question.
+Appendix A A Generalized Law of Large Numbers for Random Variables with Arbitrary Dependence
 Consider a infinitely long string ${\mathbf{Y}}=[Y_{1}~Y_{2}~\dots~Y_{n}~\dots]$ consisting of a sequence of symbols $Y_{n}\in\mathcal{Y}$ generated stochastically from a finite dictionary $\mathcal{Y}$.
@@
 We also do not assume any Markovian or conditional independence properties.
-B.1 Conditional Means and Sample Means for Sequences of Dependent Random Variables
+A.1 Conditional Means and Sample Means for Sequences of Dependent Random Variables
 Consider a sequence of random variables ${\mathbf{Z}}=[Z_{1}~Z_{2}~\dots~Z_{n}~\dots]$ whose $n$’th element $Z_{n}$ is a function of the first $n$ symbols $Y_{1},Y_{2},\dots,Y_{n}$ i.e.
@@
 We define the conditional mean and conditional variance of the random variables $Z_{n}$:
-$\mu_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)g_{n}({\mathbf{Y}}_{n})\equiv\operatorname*{{E}}\left[Z_{n}|{\mathbf{Y}}_{n-1}\right]$ (26)
-$V_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)\left(g_{n}({\mathbf{Y}}_{n})-\mu_{n}^{({\mathbf{Y}})}\right)^{2}\equiv\operatorname*{{E}}\left[Z_{n}^{2}|{\mathbf{Y}}_{n-1}\right]-\left(\mu_{n}^{({\mathbf{Y}})}\right)^{2}$ (27)
+$\mu_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)g_{n}\left([{\mathbf{Y}}_{n-1}~y]\right)\equiv\operatorname*{{E}}\left[Z_{n}|{\mathbf{Y}}_{n-1}\right]$ (28)
+$V_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)\left(g_{n}\left([{\mathbf{Y}}_{n-1}~y]\right)-\mu_{n}^{({\mathbf{Y}})}\right)^{2}\equiv\operatorname*{{E}}\left[Z_{n}^{2}|{\mathbf{Y}}_{n-1}\right]-\left(\mu_{n}^{({\mathbf{Y}})}\right)^{2}$ (29)
 where the distribution $p^{({\mathbf{Y}})}_{n}(y)\doteq\Pr(Y_{n}=y|{\mathbf{Y}}_{n-1})$.
@@
 Clearly, $\operatorname*{{E}}\left[Z_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{n}^{({\mathbf{Y}})}\right]$ and $\operatorname*{{E}}\left[S_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{S_{n}}^{({\mathbf{Y}})}\right]$.
-Lemma B.1.
+The key idea is that the sequence $Z_{m}-\mu_{m}^{({\mathbf{Y}})}$ is a martingale difference sequence, and $S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}$ is a martingale for any arbitrary distribution of the $Z_{m}$’s.
+Lemma A.1.
 The variance of the zero-mean random variable $S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}$ can be written as sums of contributions from each of the $Z_{m}$’s. Specifically, we have:
@@
 $V_{S_{l+1}}$ $=\operatorname*{{E}}\left[\left(S_{l}+Z_{l+1}-\mu_{S_{l}}^{({\mathbf{Y}})}-\mu_{l+1}^{({\mathbf{Y}})}\right)^{2}\right]$ (34)
-$=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]-2\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]$ (33)
+$=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]+2\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]$ (35)
 We will now show that the last term in (35) is zero. We apply the law of iterated expectations to write:
@@
 Remark.
-An alternative way to arrive at (35) is to treat $\mu_{l+1}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{l+1}\big{|}{\mathbf{Y}}_{l}\right]$ as an estimate of $Z_{l+1}$ given ${\mathbf{Y}}_{l}$, and then use the famous Orthogonality Principle to argue that the estimation error $\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)$ must be uncorrelated with any function of ${\mathbf{Y}}_{l}$.
+An alternative way to arrive at (37) is to treat $\mu_{l+1}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{l+1}\big{|}{\mathbf{Y}}_{l}\right]$ as an estimate of $Z_{l+1}$ given ${\mathbf{Y}}_{l}$, and then use the Orthogonality Principle to argue that the estimation error $\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)$ must be uncorrelated with any function of ${\mathbf{Y}}_{l}$.
 More formally, $\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)$ is $\mathcal{F}_{l}$-measurable and is therefore orthogonal to $\left(Z_{l+1}-\operatorname*{{E}}\left[Z_{l+1}|\mathcal{F}_{l}\right]\right)$.
@@
 where we used the induction assumption (33). We have shown that if (32) holds for $n=l$, then (32) must also hold for $n=l+1$, thus completing the induction. ∎
-Lemma B.2.
+Lemma A.2.
 The sequence $\frac{1}{n}\left(S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right)$ converges to zero in probability:
@@
 Proof.
-Using Lemma B.1, we can write:
+Using Lemma A.1, we can write:
 $V_{S_{n}}$ $=\sum_{m=1}^{n}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]\leq nG^{2}$ (40)
@@
 ∎
-Lemma B.2 states that the sample means of the two random sequences $\{Z_{n}\}$ and $\{\mu^{({\mathbf{Y}})}_{n}\}$ must be close to each other asymptotically.
+Lemma A.2 states that the sample means of the two random sequences $\{Z_{n}\}$ and $\{\mu^{({\mathbf{Y}})}_{n}\}$ must be close to each other asymptotically.
 Since we have made very few assumptions about the random variables involved, we cannot assume that either of the sequences $\{Z_{n}\}$ or $\{\mu^{({\mathbf{Y}})}_{n}\}$ themselves converge in any sense to an asymptotic limit.
@@
 a statement that the sample mean of a sequence of random variables $\{Z_{n}\}$ with arbitrary dependence converges asymptotically to a suitably defined average quantity in probability.
-Proposition B.1.
+Proposition A.1.
 A Weak Law of Large Numbers.
@@
 $\mu^{({\mathbf{Y}})}\equiv\lim\limits_{n\rightarrow\infty}\frac{1}{n}\sum\limits_{m=1}^{n}\mu_{m}^{({\mathbf{Y}})}$ exists, then the sequence $\frac{1}{n}S_{n}$ converges in probability to $\mu^{({\mathbf{Y}})}$:
-$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=1}^{n}Z_{n}$ $\to_{p}\mu^{({\mathbf{Y}})}~{}~{}\mathrm{or}~{}\lim_{n\to\infty}\Pr\left(\left|\frac{1}{n}S_{n}-\mu^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~{}\forall\epsilon>0$ (39)
+$\frac{1}{n}\sum_{m=1}^{n}Z_{n}$ $~\begin{subarray}{c}\rightarrow\\ p\end{subarray}~\mu^{({\mathbf{Y}})}~~\mathrm{or}~\lim_{n\to\infty}\Pr\left(\left|\frac{1}{n}S_{n}-\mu^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~\forall\epsilon>0$ (41)
 Proof.
-Combining Lemma B.2 with the observation $\frac{1}{n}\mu_{S_{n}}^{({\mathbf{Y}})}\rightarrow\mu^{({\mathbf{Y}})}$ completes the proof. ∎
-B.2 Discussion
-Proposition B.1 is remarkable because unlike traditional versions of the LLN, it makes no assumptions whatsoever about the dependence or correlation between the symbols $Y_{n}$.
-It does not even require for instance that the correlation between $Y_{1}$ and $Y_{n}$ should decrease as $n$ becomes large.
-The two extreme cases of strong dependence and complete independence for the symbols $Y_{n}$ turn out to be two simple special cases for our proposed LLN:
-  • If the symbols $Y_{n}$ are independent, Proposition B.1 simply reduces to the classical weak LLN.
-  • If the symbols are perfectly correlated, the $Z_{m}$ become deterministic functions of the first token, and the LLN then becomes a trivial observation.
-Appendix C Supplemental Material
-C.1 Additional Experimental Context and Results
-In this work, all experiments were ran on a high-performance computing server running Linux (Ubuntu).
-AMD EPYC 7413 24-Core Processor, 48 logical cores, 96 GB RAM, four NVIDIA A100-SXM4-80GB GPUs.
-On this setup, each run of a text generation (i.e., running GPT-2 to generate Nmax = 512 tokens, Llama-8B to generate Nmax = 1024 tokens and Nmax = 10240 tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds to minutes, with each experiment taking about 90 minutes to run in the case of 10,240 steps.
-We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.3.
-Figure 5 shows additional results similar to Fig.
-2 where we naively varied the seed value of the random number generator from $209-214$ for Llama-8B.
-We present these to show that Fig.
-2’s results are representative of the general case and were not cherry-picked.
-In all cases, $K=100$.
-(a) Seed=209.
-(b) Seed=210.
-(c) Seed=211.
-(d) Seed=212.
-(e) Seed=213.
-(f) Seed=214.
-Figure 5: Log-perplexity and empirical entropy for Llama 3.1 of strings generated by the same model with top-100 sampling.
+Combining Lemma A.2 with the observation $\frac{1}{n}\mu_{S_{n}}^{({\mathbf{Y}})}\rightarrow\mu^{({\mathbf{Y}})}$ completes the proof. ∎